### v1

# An existence and uniqueness result to evolution equations with sign-changing pseudo-differential operators and its applications to logarithmic Laplacian operators and second-order differential operators without ellipticity

Abstract.
We broaden the domain of the Fourier transform to contain all distributions without using the Paley-Wiener theorem and devise a new weak formulation built upon this extension. This formulation is applicable to evolution equations involving pseudo-differential operators, even when the signs of their symbols may vary over time. Notably, our main operator includes the logarithmic Laplacian operator $\log(-\Delta)$ and a second-order differential operator whose leading coefficients are not positive semi-definite.
Key words and phrases:
Pseudo-differential operators with sign-changing symbols, Fourier transform beyond tempered distributions, logarithmic Laplacian, Cauchy problem without ellipticity, Weighted Bessel potential spaces
2020 Mathematics Subject Classification:
35S05, 35S10, 35A01, 35D30, 47G30
This research was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No.2020R1A2C1A01003959)
1. Introduction
Pseudo-differential operators emerge as a mathematical generalization of differential operators. An illustrative form of a pseudo-differential operator is expressed as
$P(x,-i\nabla)u(x)=\mathcal{F}^{-1}\left[\sigma(x,\xi)\mathcal{F}[u](\xi)\right](x),$
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and the inverse Fourier transforms on $\mathbb{R}^{d}$, respectively. In general, pseudo-differential operators lack local properties, which make conventional perturbation methods from partial differential equation theories inapplicable. Consequently, studying the well-posedness of the equation
$P(x,-i\nabla)u(x)=f(x)$ (1.1)
within a specific framework becomes highly challenging without imposing very strong assumptions on both the symbol $\sigma(x,\xi)$ and the data $f$. However, if one considers pseudo-differential operators which are independent of the space variable $x$, there might be intriguing theories for addressing a general symbol $\sigma(\xi)$ and data $f$ within a weak formulation. Specifically, when considering a polynomial symbol $\sigma(\xi)$, there exist exceptionally elegant theories regarding the existence and uniqueness of solutions to (elliptic) equations like (1.1) in the weak sense. Remarkably, these theories hold even in the absence of strong mathematical conditions on $\sigma$ and $f$. In particular, a solution $u$ exists in the sense of distributions if the symbol $\sigma$ is a polynomial, $\sigma(\xi)\not\equiv 0$, and $f$ is a distribution with a compact support on $\mathbb{R}^{d}$. Even more surprisingly, there always exists a distribution solution $u$ to (1.1) for any distribution $f$ if the symbol $\sigma$ is a non-zero polynomial, which is shown based on the Paley-Wiener theorem and duality arguments. This existence result holds even though a nonzero polynomial symbol $\sigma(\xi)$ does not satisfy an ellipticity condition, since the set of all zeros of a non-zero polynomial has a lower dimension. However, the uniqueness of a weak solution $u$ is not guaranteed if there is no ellipticity condition. Thus ellipticity is necessary and sufficient to guarantee both existence and uniqueness of a weak solution to (1.1) even for a polynomial symbol which is independent of $x$. Moreover, this well-posedness theory can be generalized to accommodate non-polynomial symbols $\sigma$ and tempered distributions $f$ by imposing supplementary conditions involving growth, regularity, and ellipticity on the symbol $\sigma$. For more in-depth information, we recommend that the reader consult the books [10, 11, 24, 25, 27]. It is noteworthy that for a polynomial symbol $\sigma(\xi)$, the sufficient condition to ensure the weak existence of (1.1) with $f$ having a distribution with compact support is simply that $\sigma(\xi)$ is not identically zero. This condition is feasible through a translation of the polynomial $\sigma(\xi)$ and using properties of convolution and analytic functions (see, for instance, [10, Theorem 7.3.10] and [11, Theorem 10.3.1]). However, this approach may not be applicable for non-polynomial symbols, particularly in scenarios where the symbols lack regularity. In such scenarios, an ellipticity condition on $\sigma$ becomes crucial, as it prompts the consideration of the reciprocal $\frac{1}{\sigma}$ to derive a fundamental solution (Green’s function).
On the other hand, when we consider the evolutionary (or parabolic) counterparts of these theories, the reciprocal $\frac{1}{\sigma}$ could not emerge as a factor in acquiring a fundamental solution if the Fourier transforms are taken solely with respect to the space variable. Moreover, this approach to take the Fourier transforms only with respect to the spatial variable facilitates examinations of time-dependent symbols. Certainly, at least in a formal sense, a fundamental solution to the following simple second-order evolution equation
$\partial_{t}u(t,x)=a^{ij}(t)u_{x^{i}x^{j}}(t,x),\quad u(0)=\delta_{0}$
is given by
$\mathcal{F}_{\xi}^{-1}\left[\exp\left(-\int_{0}^{t}a^{ij}(s)\xi^{i}\xi^{j}\mathrm{d}s\right)\right],$ (1.2)
where $\delta_{0}$ is the Dirac delta measure centered at the origin and $\mathcal{F}_{\xi}^{-1}$ denotes the inverse Fourier transform with respect to the variable $\xi$. Similarly, we could claim that a solution $u$ to
$\partial_{t}u(t,x)=\mathcal{F}^{-1}\left[\sigma(t,\xi)\mathcal{F}[u(t,\cdot)](\xi)\right]$
$u(0,x)=\delta_{0}$ (1.3)
is given by
$u(t,x)=\mathcal{F}_{\xi}^{-1}\left[\exp\left(\int_{0}^{t}\sigma(s,\xi)\mathrm{d}s\right)\right](x).$
This leads us to ponder whether ellipticity can be eliminated in a weak formulation of evolution equations for general symbols. We found that providing a positive answer to this question is indeed valid when investigating evolution equations through a novel weak formulation. This paper is an outcome of developing a suitable weak formulation to eliminate all elliptic conditions from main operators in evolution equations even though symbols are not regular (thus definitely are not polynomials). This new weak formulation relies entirely on an extension of the Fourier and inverse Fourier transforms. Traditionally, the class of all tempered distributions has been recognized as an extensive class suitable for these transforms since both transforms become automorphisms on this class. However, there exists a remarkable method to extend the domain of the Fourier transform to include all distributions. This is possible due to the Paley-Wiener theorem. More precisely, the Fourier transform of a distribution on $\mathbb{R}^{d}$ becomes an element of the dual space of the subspace of entire functions $F$ so that for any multi-index $\alpha$ and $N>0$,
$|D^{\alpha}F(z)|\leq C(1+|z|)^{-N}\mathrm{e}^{B|\Im(z)|}\quad\forall z\in\mathbb{C}^{d},$ (1.4)
where $C$ and $A$ are positive constants depending on $\alpha$ and $N$, and $\Im(z)$ denotes the imaginary part of the complex vector $z$ (cf. [24, Proposition 4.1]). Nonetheless, this extension of the Fourier transform is not appropriate to deal with initial and inhomogeneous data on the Euclidean spaces without having analytic continuations. Thus we propose a new approach which enables the Fourier and inverse Fourier transforms to apply for all distributions without considering analytic continuations. We will provide a straightforward explanation of these extensions later in the introduction, following the presentation of the main equations. Alternatively, you can examine the comprehensive formulations with all the specifics in Definitions 2.1, 2.2, and 2.3 in the upcoming section.
Now, we succinctly present our equations, which naturally arise as a generalization of (1.3) :
$\begin{cases}\partial_{t}u(t,x)=\psi(t,-i\nabla)u(t,x)+f(t,x),\quad&(t,x)\in(0,T)\times\mathbb{R}^{d},\\ u(0,x)=u_{0},\quad&x\in\mathbb{R}^{d}.\end{cases}$ (1.5)
Here the main operator $\psi(t,-i\nabla)$ is given by
$\psi(t,-i\nabla)u(t,x):=\mathcal{F}^{-1}\left[\psi(t,\cdot)\mathcal{F}[u(t,\cdot)]\right](x)$ (1.6)
and called a time-measurable pseudo-differential operator, where $\psi(t,\xi)$ represents a complex-valued (Borel) measurable function defined on $(0,T)\times\mathbb{R}^{d}$ and is commonly referred to as a symbol. To maintain consistency with the notations used in our previous findings, we substitute $\psi(t,\xi)$ for $\sigma(t,\xi)$. This adjustment is made since $\sigma$ conventionally represents the coefficients associated with random noise terms in the realm of stochastic partial differential equations (cf. [26]).
Numerous attempts have been undertaken to ease the conditions on symbols while still ensuring the well-posedness of partial differential equations that involve these time-measurable pseudo-differential operators. It is reasonable to anticipate that the symbol may not require a regularity condition concerning the time variable $t$, as evidenced by many results pertaining to second-order parabolic equations and diffusion processes (cf. [27, 31]). However, a regularity condition on $\psi(t,\xi)$ with respect to $\xi$ appears to be essential for defining the operator $\psi(t,-i\nabla)$ in a strong space, such as an $L_{p}$-space or a Hölder space, as indicated by multiplier theories in Fourier analysis (cf. [8, 9]). One of the most famous examples is Mihlin’s multiplier and its evolutionary counterpart is given by
$|D_{\xi}^{\alpha}\psi(t,\xi)|\lesssim|\xi|^{\nu-|\alpha|}\quad\forall(t,\xi)\in(0,T)\times\mathbb{R}_{0}^{d}$ (1.7)
for all multi-index $\alpha$ so that $|\alpha|\leq\lfloor\frac{d}{2}\rfloor+1$, where $\mathbb{R}_{0}^{d}=\mathbb{R}^{d}\setminus\{0\}$, $\nu$ is a positive constant representing an order of the operator, and $\lfloor\frac{d}{2}\rfloor$ is the largest integer which is less than or equal to $\frac{d}{2}$.
Furthermore, a regularity condition on $\psi(t,\xi)$ alone is insufficient to ensure a strong solution to (1.5), as observed in the case of a fundamental solution for the second-order equation in (1.2). In essence, the term $\exp\left(-\int_{0}^{t}a^{ij}(s)\xi^{i}\xi^{j}\mathrm{d}s\right)$ in (1.2) may exhibit exponential growth if there is no elliptic condition on the coefficients $a^{ij}$. Thus, we introduce the following strong elliptic condition: there exists a positive constant $\nu>0$ such that
$-\Re[\psi(t,\xi)]\geq\nu|\xi|^{\gamma}\quad\forall\xi\in\mathbb{R}^{d},$ (1.8)
where $\Re[\psi(t,\xi)]$ represents the real part of the complex number $\psi(t,\xi)$.
Drawing upon the criteria outlined in (1.7) and (1.8), extensive researches conducted by the authors and collaborators have consistently shown robust well-posedness for the equation (1.5). Examples of such research can be found in [1, 2, 5, 15, 16, 17, 18]. These operators exhibit a close connection to non-local operators and serve as generators of stochastic processes. Additionally, for a discussion on the strong well-posedness of these equations with non-local operators and generators of stochastic processes in $L_{p}$-spaces, we direct readers to the references [3, 4, 6, 12, 13, 14, 20, 22, 29, 30, 32].
Historically, our attention has been primarily directed towards investigating the strong well-posedness in $L_{p}$ or Hölder spaces. This emphasis was rooted in a longstanding belief that the uniform ellipticity, as expressed in (1.8), was indispensable and could not be eliminated, even when considering weak solutions to equations like (1.5). The reasoning seems clear and uncomplicated. If we examine the quantity $\exp[\psi(t,\xi)]$ without an ellipticity, then it becomes apparent that this function could have exponential growth in general, hindering it from being a tempered distribution on $\mathbb{R}^{d}$. Thus, the inverse Fourier transform of the function $\exp(\psi(t,\xi))$ loses its meaningfulness, particularly when considered as a fundamental solution to (1.5), even in a weak sense if we do not consider extensions of equations to $\mathbb{C}^{d}$ after taking the Fourier transform. Moreover, it is impossible to find a nice analytic continuation of $\psi(t,\xi)$ if there is no regularity condition on $\psi(t,\xi)$ at all. Hence, it becomes crucial to ensure that the real part of the symbol $\psi(t,\xi)$ remains non-positive in order to guarantee the existence of a fundamental solution even as a tempered distribution-valued function. In other words, the strong elliptic condition outlined in (1.8) (or non-positivity of the symbol) seems to be a suitable requirement, even in the context of a weak formulation. Therefore, initially we were inclined to believe that a weak well-posedness theory for (1.5) with a strong elliptic condition (1.8) might not be appealing for research, even if regularity conditions such as (1.7) could be eliminated in the weak formulation. However, we recently discovered a method to extend the domain of the Fourier and the inverse Fourier transforms to all distributions on $\mathbb{R}^{d}$ without using the Paley-Wiener theorem. In particular, it becomes possible to assign a certain meaning to the term $\mathcal{F}^{-1}\left[\exp(\psi(t,\xi))\right]$ instead of an element in the dual space of entire functions $F$ on $\mathbb{C}^{d}$ so that (1.4) holds, even when the sign of the real part of the symbol $\psi(t,\xi)$ is strictly positive. With these extensions, we ultimately achieve a satisfactory weak well-posedness result for (1.5) without imposing any ellipticity and regularity conditions. A certain local integrability on the symbol $\psi(t,\xi)$ is sufficient to establish weak well-posedness of (1.5). We emphasize that not only the existence, but also the uniqueness of a weak solution is obtained even though a weak solution to equation (1.1) is not unique even for a polynomial symbol $\sigma(\xi)$ without ellipticity. For further details, see Theorem 2.21 below.
Next, we offer a brief overview of a fundamental idea behind our weak formulation. It is essential to recognize that trying to create a weak formulation across the entire domain $(0,T)\times\mathbb{R}^{d}$ is unfeasible due to the fact that a symbol $\psi(t,\xi)$ varies with respect to the time. Additionally, $(0,T)$ could be a finite interval where Schwartz’s functions do not work effectively. As a result, our testing actions for the equations are limited to only the spatial variables. The core aspect of our approach lies in utilizing a different class of test functions. Specifically, we use $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ as a class of test functions instead of $\mathcal{D}(\mathbb{R}^{d})$. Here, $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ represents a subset of Schwartz functions whose Fourier transforms are belonging to the class $\mathcal{D}(\mathbb{R}^{d})$ and $\mathcal{D}(\mathbb{R}^{d})$ is the class of all functions that are infinitely differentiable and have compact supports. We establish the Fourier transform and its inverse for a distribution $v$ defined on $\mathbb{R}^{d}$ as linear functionals acting on $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$. This is achieved by applying straightforward operations on $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, which are motivated by Plancherel’s theorem. In other words, we define
$\langle\mathcal{F}^{-1}[v],\varphi\rangle=\langle v,\mathcal{F}[\varphi]\rangle\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$ (1.9)
and
$\langle\mathcal{F}[v],\varphi\rangle=\langle v,\mathcal{F}^{-1}[\varphi]\rangle\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
Specifically, we have the option to consider the Fourier transform and its inverse for any function which is locally integrable.
Now, suppose that we have a solution $u$ to equation (1.5) such that the Fourier transform of $u$ has a realization on $\mathbb{R}^{d}$. Furthermore, we also assume that the product of $\psi(t,\xi)$ and the Fourier transform of $u$ with respect to $\xi$ is locally integrable. Then, the function referred to in (1.6) becomes meaningful due to the revised definition of the inverse Fourier transform presented in (1.9).
To conclude, we can define $u$ as a weak solution to (1.5) if, for every $\varphi$ belonging to the space $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, the following equation holds:
$\left\langle u(t,\cdot),\varphi\right\rangle=\langle u_{0},\varphi\rangle+\int_{0}^{t}\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\rangle\mathrm{d}s+\int_{0}^{t}\left\langle f(s,\cdot),\varphi\right\rangle\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$
This straightforward concept enables us to naturally manage the exponential growth that may arise when the real part of the symbol $\psi(t,\xi)$ is sign-changing.
It is worth noting that we do not rely on any arguments related to weak pre-compactness to derive our main theorems. Instead, our results are solely grounded in elementary techniques, such as Hölder and Minkowski inequalities, combined with well-known properties of the Fourier and inverse Fourier transforms.
This paper comprises seven sections. Section 1 serves as an introductory section, presenting an overview of the results. Our primary theorems are introduced in Section 2. In Section 3, we revisit the classical Bessel potential spaces and extend them to incorporate weighted variations. Section 4 is dedicated to establishing the uniqueness of a solution $u$ for (1.5) under a slightly relaxed condition. The proof of the main theorem is provided in Section 5, along with noteworthy corollaries for practical applications. Section 6 focuses on handling logarithmic operators as a specific application of our theorem, while Section 7 addresses second-order differential operators without ellipticity.
We finish this section with the notations used in the article.
  • Let $\mathbb{N}$, $\mathbb{Z}$, $\mathbb{R}$, $\mathbb{C}$ denote the natural number system, the integer number system, the real number system, and the complex number system, respectively. For $d\in\mathbb{N}$, $\mathbb{R}^{d}$ denotes the $d$-dimensional Euclidean space.
  • For $i=1,...,d$, a multi-index $\alpha=(\alpha_{1},...,\alpha_{d})$ with $\alpha_{i}\in\{0,1,2,...\}$, and a function $g$, we set
    $\frac{\partial g}{\partial x^{i}}=D_{x^{i}}g,\quad D^{\alpha}g=D_{x^{1}}^{\alpha_{1}}\cdot...\cdot D^{\alpha_{d}}_{x^{d}}g,\quad|\alpha|:=\sum_{i=1}^{d}\alpha_{i}.$
    For $\alpha_{i}=0$, we define $D^{\alpha_{i}}_{x^{i}}g=g$. If $g=g(t,x):\mathbb{R}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, then we denote
    $D_{x}^{\alpha}g=D_{x^{1}}^{\alpha_{1}}\cdot...\cdot D^{\alpha_{d}}_{x^{d}}g,$
    where $\alpha=(\alpha_{1},\cdots,\alpha_{d})$.
  • We employ the identical symbol “$\sup$” to represent both the supremum and the essential supremum by slightly abusing the notation $\sup$. On occasion, “$\operatorname*{ess\,sup}$” may be utilized to emphasize its role as the essential supremum.
  • We use $\mathcal{D}(\mathbb{R}^{d})$ to denote the space of all infinitely differentiable functions with compact supports. $\mathcal{S}(\mathbb{R}^{d})$ represents the Schwartz space on $\mathbb{R}^{d}$ and the topology on $\mathcal{S}(\mathbb{R}^{d})$ is generated by the Schwartz semi-norms $\sup_{x\in\mathbb{R}^{d}}|x^{\alpha}(D^{\beta}f)(x)|$ for all multi-indexes $\alpha$ and $\beta$. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is used to denote the dual space of $\mathcal{S}(\mathbb{R}^{d})$, i.e. $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is the space of all tempered distributions on $\mathbb{R}^{d}$. Additionally, we assume that $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ is a topological space equipped with the weak*-topology if there is no special remark about the topology on $\mathcal{S}^{\prime}(\mathbb{R}^{d})$.
  • Let $F$ be a normed space and $(X,\mathcal{M},\mu)$ be a measure space.
    • $\mathcal{M}^{\mu}$ denotes the completion of $\mathcal{M}$ with respect to the measure $\mu$.
    • For $p\in[1,\infty)$, the space of all $\mathcal{M}^{\mu}$-measurable functions $f:X\to F$ with the norm
      $\left\|f\right\|_{L_{p}(X,\mathcal{M},\mu;F)}:=\left(\int_{X}\left\|f(x)\right\|_{F}^{p}\mu(\mathrm{d}x)\right)^{1/p}<\infty$
      is denoted by $L_{p}(X,\mathcal{M},\mu;F)$. We also denote by $L_{\infty}(X,\mathcal{M},\mu;F)$ the space of all $\mathcal{M}^{\mu}$-measurable functions $f:X\to F$ with the norm
      $\|f\|_{L_{\infty}(X,\mathcal{M},\mu;F)}:=\inf\left\{r\geq 0:\mu(\{x\in X:\|f(x)\|_{F}\geq r\})=0\right\}<\infty.$
      We usually omit the given measure or $\sigma$-algebra in the notations of $L_{p}$-spaces if there is no confusion (e.g. Lebesgue (or Borel) measure and $\sigma$-algebra). We similarly leave out the representation of $F$ when it takes on the scalar values such as $\mathbb{R}$ or $\mathbb{C}$.
  • Let $f(t)$ and $g(t)$ be Lebesgue measurable (or Borel measurable) functions on $\mathcal{U}$. We write
    $f(t)=g(t)~{}(a.e.)~{}t\in\mathcal{U}$
[…]
  • For $R>0$,
    $B_{R}(x):=\{y\in\mathbb{R}^{d}:|x-y|<R\},\quad\overline{B_{R}(x)}:=\{y\in\mathbb{R}^{d}:|x-y|\leq R\}$
  • For a measurable function $f$ on $\mathbb{R}^{d}$, we denote the $d$-dimensional Fourier transform of $f$ by
    $\mathcal{F}[f](\xi):=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{-i\xi\cdot x}f(x)\mathrm{d}x$
    and the $d$-dimensional inverse Fourier transform of $f$ by
    $\mathcal{F}^{-1}[f](x):=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}f(\xi)\mathrm{d}\xi.$
    Moreover, for a function $f(t,x)$ defined on $(0,T)\times\mathbb{R}^{d}$, we use the notation
    $\mathcal{F}_{x}[f](\xi):=\mathcal{F}[f(t,x)](\xi):=\mathcal{F}[f(t,\cdot)](\xi)=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{-i\xi\cdot x}f(t,x)\mathrm{d}x$
    and it is called the Fourier transform of $f$ with respect to the space variable. On the other hand, for the inverse Fourier transform of $f(t,\xi)$ with respect to the space variable, we use the notation
    $\mathcal{F}_{\xi}^{-1}[f](x):=\mathcal{F}^{-1}[f(t,\xi)](x):=\mathcal{F}^{-1}[f(t,\cdot)](x)=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}f(t,\xi)\mathrm{d}\xi.$
    For the sake of simplicity, we often omit the subscripts denoted by $x$ and $\xi$. It is well-known that both the Fourier transform and the inverse Fourier transform have extensions on $L_{1}(\mathbb{R}^{d})+L_{2}(\mathbb{R}^{d})$. By slightly abusing the notation, we use the same notation $\mathcal{F}$ and $\mathcal{F}^{-1}$ to denote these extensions. We discuss more details some properties of these extensions and corresponding inversion theorems in Section 3.
  • We write $\alpha\lesssim\beta$ if there is a positive constant $N$ such that $\alpha\leq N\beta$. We use $\alpha\approx\beta$ if $\alpha\lesssim\beta$ and $\beta\lesssim\alpha$. In particular, we use the notation $\alpha\lesssim_{a,b,\cdots}\beta$ if the constant $N$ so that $\alpha\leq N\beta$ depends only on $a,b,\cdots$. Moreover, if we write $N=N(a,b,\cdots)$, this means that the constant $N$ depends only on $a,b,\cdots$. A generic constant $N$ may change from a location to a location, even within a line. The dependence of generic constants is usually specified in each statement of theorems, propositions, lemmas, and corollaries.
  • For $z\in\mathbb{C}$, $\Re[z]$ denotes the real part of $z$, $\Im[z]$ is the imaginary part of $z$ and $\bar{z}$ is the complex conjugate of $z$.
2. Settings and main results
[…]
$\psi(t,-i\nabla)u(t,x):=\mathcal{F}^{-1}\left[\psi(t,\cdot)\mathcal{F}[u(t,\cdot)]\right](x)$ (2.1)
in the whole paper. All the Fourier and the inverse Fourier transforms in the paper are taken only with respect to the space variable and thus they are considered only on $\mathbb{R}^{d}$ even in a weak sense.
Our main goal of this paper is to show the existence and uniqueness of a weak solution $u$ to the Cauchy problem (1.5). Note that it is not easy to recover a strong solution to (1.5) even in $L_{p}$-spaces if there are no strong mathematical conditions on the symbol and data. A naive condition is that for each $t\in(0,T)$, $\psi(t,\cdot)\mathcal{F}[u](t,\cdot)\in L_{p}(\mathbb{R}^{d})$ with $p\in[1,2]$. Then the function $t\mapsto\psi(t,-i\nabla)u(t,\cdot)$ is well-defined as an $L_{p^{\prime}}(\mathbb{R}^{d})$-valued function defined on $(0,T)$ due to the Riesz-Thorin Theorem, where $p^{\prime}$ denotes the Hölder conjugate of $p$, i.e.
$p^{\prime}=\begin{cases}&\frac{p}{p-1}\quad\text{if $p\in(1,2]$}\\ &\infty\quad\text{if $p=1$}.\end{cases}$
We revisit Riesz-Thorin’s Theorem in Section 3, providing more specific details about the constants involved.
On the other hand, it appears that there is no need for numerous mathematical conditions when approaching the operator described in (2.1) in a weak sense. In particular, if $\psi(t,\cdot)\mathcal{F}[u](t,\cdot)\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, then $t\mapsto\psi(t,-i\nabla)u(t,\cdot)$ is well-defined as an $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued function defined on $(0,T)$, where $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ denotes the class of all tempered distributions on $\mathbb{R}^{d}$. Nevertheless, if the function $\psi(t,\cdot)\mathcal{F}[u]$ is only locally integrable, it does not qualify as a tempered distribution on $\mathbb{R}^{d}$. Consequently, the operator in (2.1) is not well-defined in this scenario, as it may have exponential growth. However, local integrability of $\psi(t,\cdot)\mathcal{F}[u]$ is sufficient to establish it as a distribution on $\mathbb{R}^{d}$. Hence, if we suggest a new technique for expanding the domain of the inverse Fourier transform to encompass all distributions, then the operator defined in (2.1) could be meaningful even with only local integrability on $\psi(t,\cdot)\mathcal{F}[u]$. This extension is especially important when dealing with functions that exhibit exponential growth. Additionally, it is well-known that there exists an elegant method to define the Fourier transform of a distribution on $\mathbb{R}^{d}$ as an element in the dual space of a subspace of entire functions $F$ on $\mathbb{C}^{d}$ so that (1.4) holds. However, this extension is not appropriate to fulfill our well-posedness theory since our symbol $\sigma(t,\xi)$ is not a polynomial type in general. In other words, we cannot consider an analytic continuation of $\sigma(t,\xi)$ on $\mathbb{C}^{d}$ for each $t$. Thus, we suggest a new extension of the Fourier transform to include all distributions below in Definitions 2.2 and 2.3. Remarkably, we have not come across a suitable reference discussing this type extension of the Fourier transform beyond tempered distributions without using analytic continuations to the best of our knowledge.
Utilization of distribution theories has significantly advanced our understanding of partial differential equations on a broader scale, as noted in [7, 10, 11]. Nevertheless, a conventional weak formulation utilizing distributions appears ineffective in the absence of strong ellipticity on the symbols for our evolution equations due to weak mathematical conditions with respect to the time variable $t$ and possibility having exponential growth as highlighted in the introduction. More specifically, if we allow the symbol $\psi(t,\xi)$ to change its sign, then our equation (1.5) includes second-order evolution equations without the parabolicity, which clearly shows that classical weak formulations taken for both time and space variables simultaneously, cannot be applicable since the coefficients are depending on the time, $(0,T)$ could be finite interval, and data $u_{0}$ and $f$ might have not compact supports in general. Moreover, exponential growth could emerge naturally as a result of the influence of exponential functions generated by time evolution when examining evolution equations instead of stationary equations if there is no ellipticity. This occurrence is particularly evident when Fourier transforms are exclusively applied to the spatial variable to accommodate variations in symbols with respect to the time variable.
Thus, it is reasonable to anticipate that the standard framework of weak formulations typically tested with $\mathcal{D}(\mathbb{R}^{d})$ or $\mathcal{S}(\mathbb{R}^{d})$ is insufficient to address well-posedness of our equation (1.5). A typical form of a solution to (1.5) clarifies these reasons. Indeed, at least in a formal sense, a solution $u$ to (1.5) can be expressed as follows:
$u(t,x)=\mathcal{F}^{-1}\left[\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]\right](x)+\mathcal{F}^{-1}\left[\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s\right](x).$ (2.2)
Hence, if our symbol $\psi(t,\xi)$ shows a sign-changing behavior, it can lead to exponential growth in the frequency of our solution $u$. Nonetheless, it is widely recognized that tempered distributions do not offer adequate control over the exponential growth, as has been emphasized multiple times previously. Thus, it becomes necessary to find a more extensive class than $\mathcal{S}^{\prime}(\mathbb{R}^{d})$ to ensure well-posedness of our evolution equation (1.5), particularly when we permit symbols to change signs. However, using the class of all distributions as a direct replacement is not suitable because the Fourier transforms of distributions are not universally defined if we do not consider them as elements in a dual space of some entire functions. Therefore, we propose a new class referred to as Fourier transforms of distributions.
Definition 2.1 (Fourier transforms of distributions).
We use $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ to denote the subclass of the Schwartz class whose Fourier transform is in $\mathcal{D}(\mathbb{R}^{d})$, i.e.
$\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}):=\{\varphi\in\mathcal{S}(\mathbb{R}^{d}):\mathcal{F}[\varphi]\in\mathcal{D}(\mathbb{R}^{d})\},$
where $\mathcal{S}(\mathbb{R}^{d})$ denotes the (complex-valued) Schwartz class on $\mathbb{R}^{d}$ and $\mathcal{D}(\mathbb{R}^{d})$ denotes the class of all complex-valued infinitely differentiable functions with compact supports defined on $\mathbb{R}^{d}$. Recall that there is a topology on $\mathcal{S}(\mathbb{R}^{d})$ generated by the Schwartz semi-norms. Since $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\subset\mathcal{S}(\mathbb{R}^{d})$, there is a subspace topology on $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$. Therefore, one can consider its dual space. We use the notation $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ to denote the dual space of $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, i.e. $u\in\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ if and only if $u$ is a continuous linear functional defined on $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$. For $u\in\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ and $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, we write
$\langle u,\varphi\rangle:=u(\varphi).$
In other words, $\langle u,\varphi\rangle$ denotes the image of $\varphi$ under $u$. Defining a topology for $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ is a straightforward task. Specifically, one can adopt the weak*-topology for $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ and work with its associated Borel sets to mention measurability of functions whose range are in $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$.
Definition 2.2 (Fourier and inverse Fourier transforms of $u\in\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$).
For $u\in\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$, one can define the Fourier and inverse Fourier transforms of $u$ in the following way:
$\langle\mathcal{F}[u],\varphi\rangle:=\langle u,\mathcal{F}^{-1}[\varphi]\rangle$
and
$\langle\mathcal{F}^{-1}[u],\varphi\rangle:=\langle u,\mathcal{F}[\varphi]\rangle$
for all $\varphi\in\mathcal{D}(\mathbb{R}^{d})$. Thus for any $u\in\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$, we have $\mathcal{F}[u]\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$ and $\mathcal{F}^{-1}[u]\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$ since it is obvious that both $\mathcal{F}^{-1}[\varphi]$ and $\mathcal{F}[\varphi]$ are in $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ for all $\varphi\in\mathcal{D}(\mathbb{R}^{d})$ due to well-known properties of the Fourier and inverse Fourier transforms of a Schwartz function. The continuity of $\mathcal{F}[u]$ and $\mathcal{F}^{-1}[u]$ on $\mathcal{D}(\mathbb{R}^{d})$ to be distributions is easily obtained from properties of the Fourier transform and the Schwartz functions as well.
It is essential to reiterate that, for any $u\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$, the conventional definition of its Fourier transform or inverse Fourier transform is not applicable without considering analytic continuations as discussed previously. Nonetheless, we can ultimately establish a meaningful definition for the Fourier and inverse Fourier transforms of all distributions, treating them as elements in $(\mathcal{F}^{-1}\mathcal{D})^{\prime}(\mathbb{R}^{d})$ without the Paley-Wiener theorem.
Definition 2.3 (Fourier and inverse Fourier transforms of $u\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$).
Let $u\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$. We define the Fourier and inverse Fourier transforms of $u$ as follows:
$\langle\mathcal{F}[u],\varphi\rangle:=\langle u,\mathcal{F}^{-1}[\varphi]\rangle\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$
and
$\langle\mathcal{F}^{-1}[u],\varphi\rangle:=\langle u,\mathcal{F}[\varphi]\rangle\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
It is obvious that both transforms above are well-defined since $\mathcal{F}[\varphi]$ and $\mathcal{F}^{-1}[\varphi]$ are in $\mathcal{D}(\mathbb{R}^{d})$ for all $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$.
Due to Definitions 2.2 and 2.3, any distribution $u$ on $\mathbb{R}^{d}$ becomes the Fourier transform (or inverse Fourier transform) of an element in $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ and any element in $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ is the Fourier transform (or inverse Fourier transform) of a distribution on $\mathbb{R}^{d}$. In this sense, we could say that $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$ is the space of all Fourier transforms or all inverse Fourier transforms of distributions on $\mathbb{R}^{d}$. Thus we use the notation $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ or $\mathcal{F}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ instead of $\left(\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\right)^{\prime}$.
Besides, recall that both Fourier and inverse Fourier transforms are homeomorphisms on the class of all tempered distributions with respect to weak*-topologies. Since $\mathcal{D}(\mathbb{R}^{d})\subset\mathcal{S}(\mathbb{R}^{d})$ and $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})\subset\mathcal{S}(\mathbb{R}^{d})$, it is obvious that both spaces $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ and $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ are lager than $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. We state below that both transforms still become homeomorphisms with respect to weak*-topologies. We do not give a detailed proof since the proof is almost identical with that of the tempered distributions based on properties of transforms for functions in the Schwartz class (cf. [8, Section 2.2.3] and [10, Theorem 7.1.10]).
Theorem 2.4.
  1. (i)
    Both Fourier and inverse Fourier transforms are homeomorphisms from $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ onto $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ with respect to weak*-topologies.
  2. (ii)
    Both Fourier and inverse Fourier transforms are homeomorphisms from $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ onto $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ with respect to weak*-topologies.
In total, both transforms $\mathcal{F}$ and $\mathcal{F}^{-1}$ are mappings from $\mathcal{D}^{\prime}(\mathbb{R}^{d})\cup\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ onto $\mathcal{D}^{\prime}(\mathbb{R}^{d})\cup\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$.
Therefore, we could freely apply the Fourier inversion formula to all distributions and Fourier transforms of distributions on $\mathbb{R}^{d}$ due to the above theorem. In particular, $\mathcal{F}[u]=\mathcal{F}[v]$ implies $u=v$ for all distributions $u$ and $v$ on $\mathbb{R}^{d}$.
It is crucial to grasp how distributions are manifested in practice. In our main theorem, the mathematical conditions are given by realizations of Fourier transforms of distributions. Recall that any locally integrable function $v$ on $\mathbb{R}^{d}$ is a distribution on $\mathbb{R}^{d}$ due to the identification $v$ with the mapping
$\varphi\in\mathcal{D}(\mathbb{R}^{d})\mapsto\int_{\mathbb{R}^{d}}v(x)\overline{\varphi(x)}\mathrm{d}x.$
Here $v$ is locally integrable on $\mathbb{R}^{d}$ if and only if
$\int_{|\xi|<R}v(\xi)\mathrm{d}\xi<\infty\quad\forall R\in(0,\infty),$
where $B_{R}$ denotes the Euclidean ball whose center is zero and radius is $R$, i.e.
$B_{R}:=\{y\in\mathbb{R}^{d}:|y|<R\}.$
We write $v\in L_{1,\ell oc}(\mathbb{R}^{d})$ if $v$ is locally integrable on $\mathbb{R}^{d}$. Especially, if there exists a $v\in L_{1,\ell oc}(\mathbb{R}^{d})$ such that
$\langle\mathcal{F}[u],\varphi\rangle=\left(v,\varphi\right)_{L_{2}(\mathbb{R}^{d})}:=\int_{\mathbb{R}^{d}}v(\xi)\overline{\varphi(\xi)}\mathrm{d}\xi\quad\forall\varphi\in\mathcal{D}(\mathbb{R}^{d}),$
[…]
$\langle u,\varphi\rangle:=\int_{\mathbb{R}^{d}}v(\xi)\overline{\varphi(\xi)}\mathrm{d}\xi\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
[…]
$\left\langle u,\varphi\right\rangle=\left\langle\mathcal{F}[u],\mathcal{F}[\varphi]\right\rangle=\int_{\mathbb{R}^{d}}\mathcal{F}[u](\xi)\overline{\mathcal{F}[\varphi](\xi)}\mathrm{d}\xi\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
Since $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\subset\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, there is the subspace topology on $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and the Borel sets generated by this topology. Additionally, this identification with a locally integrable function could be easily performed in any domain of $\mathbb{R}^{d}$ due to the localizations of distributions. More precisely, if there exist a open set $\mathcal{U}\subset\mathbb{R}^{d}$ and a locally integrable function $v$ on $\mathcal{U}$ so that
$\langle\mathcal{F}[u],\varphi\rangle=\int_{\mathbb{R}^{d}}v(\xi)\overline{\varphi(\xi)}\mathrm{d}\xi\quad\forall\varphi\in\mathcal{D}(\mathcal{U}),$ (2.3)
then we identify $\mathcal{F}[u]$ with $v$ on $\mathcal{U}$, where $\mathcal{D}\left(\mathcal{U}\right)$ denotes the set of all infinitely differential functions on $\mathbb{R}^{d}$ with compact supports in $\mathcal{U}$. In such a case, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ and consider $\mathcal{F}[u]$ as a locally integrable function on $\mathcal{U}$. In other words, for any $u$ in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, we say that $\mathcal{F}[u]$ has a realization $v$ on $\mathcal{U}$ or simply $\mathcal{F}[u]$ is locally integrable on $\mathcal{U}$ if (2.3) holds.
According to these realizations of elements in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, we are ready to define the operator $\psi(t,-i\nabla)$. To define the operator universally for all $t$, we slightly abuse the notation by considering $\psi(-i\nabla)$ with a function $\psi(\xi)$ on $\mathbb{R}^{d}$ below.
Definition 2.5.
Let $g\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ so that $\mathcal{F}[g]$ has a realization on an open set $\mathcal{U}$ covering the support of $\psi$ and $\xi\in\mathbb{R}^{d}\mapsto\psi(\xi)\mathcal{F}[g](\xi)\in\mathbb{C}$ is a locally integrable function on $\mathbb{R}^{d}$. Then we define $\psi(-i\nabla)g$ as an element in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ so that
$\langle\psi(-i\nabla)g,\varphi\rangle=\int_{\mathbb{R}^{d}}\psi(\xi)\mathcal{F}[g](\xi)\mathcal{F}[\varphi](\xi)\mathrm{d}\xi\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d}).$
Due to the definition provided above, equation (1.5) can be considered valid even without imposing any regularity requirement on the symbol. This will be further elaborated upon in the subsequent discussion regarding the definition of a solution, as outlined with more depth in Definition 2.10.
Next we consider classes comprising Borel measurable functions with values in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, defined almost everywhere on the interval $(0,T)$.
Definition 2.6.
We use the notation $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ to represent the set of all Borel measurable functions $f$ with values in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ defined almost everywhere on the interval $(0,T)$. In particular, $f\in L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ implies that
$\langle f(t,\cdot),\varphi\rangle:=\langle f(t),\varphi\rangle<\infty\quad(a.e.)~{}t\in(0,T)\quad\text{and}\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
Definition 2.7.
[…]
$\int_{0}^{t}\langle f(s,\cdot),\varphi\rangle\mathrm{d}s:=\int_{0}^{t}\langle f(s),\varphi\rangle\mathrm{d}s<\infty\quad\forall t\in(0,T)\quad\text{and}\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$
It is obvious that $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)\subset L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. Naturally, we could consider subspaces of $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ and $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ with values in $\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$.
Definition 2.8.
[…]
$\int_{B_{R}}|\mathcal{F}[f(t,\cdot)](\xi)|\mathrm{d}\xi:=\int_{B_{R}}|\mathcal{F}[f(t)](\xi)|\mathrm{d}\xi<\infty\quad(a.e.)~{}t\in(0,T)\quad\text{and}\quad\forall R\in(0,\infty).$
Definition 2.9.
[…]
$\int_{0}^{t}\int_{B_{R}}|\mathcal{F}[f(s,\cdot)](\xi)|\mathrm{d}\xi\mathrm{d}s:=\int_{0}^{t}\int_{B_{R}}|\mathcal{F}[f(s)](\xi)|\mathrm{d}\xi\mathrm{d}s<\infty\quad\forall(t,R)\in(0,T)\times(0,\infty).$
Generally, $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$ cannot be identified with a complex-valued function defined on $(0,T)\times\mathbb{R}^{d}$ in the typical almost everywhere sense. Thus the terms $\mathcal{F}[f(s,\cdot)]$ and $\mathcal{F}[f(s,\cdot)](\xi)$ should be understood as the notation to denote $\mathcal{F}[f(s)]$ and $\mathcal{F}[f(s)](\xi)$, respectively. On the other hand, if $f$ is a nice function defined on $(0,T)\times\mathbb{R}^{d}$, then one can understand $\mathcal{F}[f(s)](\xi)$ as
$\mathcal{F}[f(s,\cdot)](\xi):=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{-ix\cdot\xi}f(s,x)\mathrm{d}x.$
We use both notations $\mathcal{F}[f(s)](\xi)$ and $\mathcal{F}[f(s,\cdot)](\xi)$ to give the same meaning in either way.
Now assume that $u_{0}\in L_{1,\ell oc}(\mathbb{R}^{d})$ and $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$. In such cases, all the components on the right side of equation (2.2) are well-defined and meaningful, provided that the symbol $\psi(t,\xi)$ is locally bounded with the help of the extension of the inverse Fourier transform.
Finally, we suggest a definition of a weak solution to equation (1.5). We call it a Fourier-space weak solution to (1.5).
Definition 2.10 (Fourier-space weak solution).
Let $u_{0}\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, and $u\in L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. We say that $u$ is a Fourier-space weak solution to equation (1.5) if for any $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$,
$\left\langle u(t,\cdot),\varphi\right\rangle=\langle u_{0},\varphi\rangle+\int_{0}^{t}\left\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\right\rangle\mathrm{d}s+\int_{0}^{t}\left\langle f(s,\cdot),\varphi\right\rangle\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$ (2.4)
The term “Fourier-space weak solution” as described above is evidently derived from the fact that the Fourier transform is applied to the spatial variable of a solution $u$.
Remark 2.11.
Note that the term $\int_{0}^{t}\left\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\right\rangle\mathrm{d}s$ is not well-defined in general even though we consider the extended Fourier and inverse Fourier transforms. It is because the multiplication action
$\langle\psi(t,\cdot)\mathcal{F}[u](t,\cdot),\varphi\rangle:=\langle\mathcal{F}[u](t,\cdot),\psi(t,\cdot)\varphi\rangle$
is not well-defined since the symbol $\psi(t,\xi)$ does not satisfy a regularity condition in general. The easiest way to make the term well-defined is to give a sufficient smooth assumption on the symbol $\psi(t,\xi)$. However, it is not the case we are interested in our weak formulation by recalling the goal that we want to remove all regularity conditions on the symbol $\psi(t,\xi)$. Fortunately, there is a method to make the term well-defined without any regularity condition on $\psi(t,\xi)$ by using a realization of $\mathcal{F}[u(t,\cdot)]$ as discussed in Definition 2.5, which will be revisited in the next remark.
Remark 2.12.
Let $u$ be a Fourier-space weak solution to (1.5). Assume that for each $t\in(0,T)$, the function $\psi(t,\xi)\mathcal{F}[u(t,\cdot)](\xi)$ is locally integrable with respect to $\xi$. In this case, it is clear that $\psi(t,\xi)\mathcal{F}[u(t,\cdot)](\xi)$ belongs to the space of distributions $\mathcal{D}^{\prime}(\mathbb{R}^{d})$. Consequently, we can proceed to consider the inverse Fourier transform of $\psi(t,\xi)\mathcal{F}[u(t,\cdot)](\xi)$ due to Definition 2.3. This implies that the main operator component, $\psi(t,i\nabla)u(t,\cdot)$, becomes meaningful as a function with values in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$. Additionally according to definitions of actions on $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, we have
$\int_{0}^{t}\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\rangle\mathrm{d}s=\int_{0}^{t}\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s$ (2.5)
for all $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, where $(\cdot,\cdot)_{L_{2}(\mathbb{R}^{d})}$ denotes the $L_{2}(\mathbb{R}^{d})$-inner product, i.e.,
$\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}=\int_{\mathbb{R}^{d}}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\overline{\mathcal{F}[\varphi](\xi)}\mathrm{d}\xi.$
Therefore, we can use the identity in (2.5) to define $\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\rangle$ as $\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}$. In other words, the term $\psi(s,-i\nabla)g$ could be defined for any $g$ whose Fourier transform has a realization so that $\psi(s,\xi)\mathcal{F}[g](\xi)$ is locally integrable even though the symbol $\psi(t,\xi)$ is not regular at all. It seems obvious that this realization of $g$ is sufficient to exist within the support of the symbol $\psi$ which will be discussed further in the subsequent remark.
Remark 2.13.
Let $u\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. Then even the term
$\int_{0}^{t}\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s=\int_{0}^{t}\left(\mathcal{F}[u(s,\cdot)],\overline{\psi}(s,\cdot)\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s$ (2.6)
is not well-defined in general. Moreover, there exists the possibility of finding a function $u$ that does not belong to the space $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$, yet under the influence of $\psi(s,\cdot)$, the expression (2.6) can still be made meaningful. To address this issue, we aim to find appropriate functions, which will be presented in our main theorem and belong to spaces whose Fourier transforms are in weighted local $L_{p,q}$-spaces which are subspaces of $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. In particular, if $u$ is a Fourier-space weak solution to (1.5) so that the term in (2.6) is well-defined, then at least, it implies that for each $s\in(0,T)$, there exists a realization of $\mathcal{F}[u(s,\cdot)](\xi)$ on the support of $\psi(s,\xi)$. In other words, for each $s$, there exists a locally integrable function $v$ on $\mathbb{R}^{d}$ and a neighborhood $\mathcal{U}$ of $\operatorname{supp}\psi(s,\cdot)$ so that
$\langle\mathcal{F}[u(s,\cdot)],\varphi\rangle=\int_{\mathbb{R}^{d}}v(\xi)\varphi(\xi)\mathrm{d}\xi\quad\forall\varphi\in\mathcal{D}\left(\mathcal{U}\right),$
where $\operatorname{supp}\psi(s,\cdot)=\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$, $\overline{\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$ is the closure of ${\{x\in\mathbb{R}^{d}:\psi(s,x)\neq 0\}}$. However, we can give any value to $\mathcal{F}[u(s,\cdot)]$ outside of $\operatorname{supp}\psi(s,\cdot)$ since it does not influence on the terms in (2.6). Thus it might suggest that the space $L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ is not appropriate to guarantee the uniqueness of a solution $u$. By the way, there is a typical action to extend by putting
$\langle\mathcal{F}[u(s,\cdot)],\varphi\rangle=0\quad\forall\varphi\in\mathcal{D}\left(\mathbb{R}^{d}\setminus\operatorname{supp}\psi(s,\cdot)\right).$
Then the support of the distribution $\mathcal{F}[u(s,\cdot)]$ becomes a subset of $supp\psi(s,\cdot)$ and the realization $v$ satisfies $v=v\cdot 1_{supp\psi(s,\cdot)}$, where $1_{supp\psi(s,\cdot)}$ denotes the indicator function on the set $supp\psi(s,\cdot)$, i.e.
$1_{supp\psi(s,\cdot)}(x)=\begin{cases}&1\quad x\in supp\psi(s,\cdot)\\ &0\quad x\in\mathbb{R}^{d}\setminus supp\psi(s,\cdot).\end{cases}$
Remark 2.14.
In the classical weak formulation of partial differential equations, the commonly preferred class of test functions is $\mathcal{D}(\mathbb{R}^{d})$. This preference is shaped by several factors. One of the primary reasons is that $\mathcal{D}(\mathbb{R}^{d})$ serves as a well-behaved dense subset. Specifically, $\mathcal{D}(\mathbb{R}^{d})$ is dense in $L_{p}(\mathbb{R}^{d})$ for any $p\in[1,\infty)$, and it is also dense in $\mathcal{S}(\mathbb{R}^{d})$. These properties of denseness are significant considerations in favor of utilizing $\mathcal{D}(\mathbb{R}^{d})$ as a space of test functions. They play a crucial role in facilitating a connection between strong and weak solutions.
It is worth noting that the class $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ serves as an excellent alternative for a class of test functions, as it possesses the same dense properties. These properties can be easily shown based on well-known properties of $\mathcal{D}(\mathbb{R}^{d})$ and $\mathcal{S}(\mathbb{R}^{d})$. Here is a brief description of a proof. Recall that $\mathcal{D}(\mathbb{R}^{d})$ is dense in $\mathcal{S}(\mathbb{R}^{d})$ concerning the topology generated by the Schwartz semi-norms. Additionally, the Fourier transform acts as a homeomorphism from $\mathcal{S}(\mathbb{R}^{d})$ to itself. Consequently, it is almost straightforward to prove that $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ is dense in $\mathcal{S}(\mathbb{R}^{d})$. An interesting observation is that both $\mathcal{D}(\mathbb{R}^{d})$ and $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ are dense in $\mathcal{S}(\mathbb{R}^{d})$, but their intersection contains only the zero element.
We classify our solutions and data by examining their Fourier transforms. Specifically, we focus on an initial data $u_{0}$ and an inhomogeneous data $f$ in equation (1.5) whose Fourier transforms belong to weighted $L_{p}$-spaces. To establish clear definitions for these data spaces, we need to review notations associated with $L_{p}$-spaces related to weighted Lebesgue measures. It is important to note that our weighted $L_{p}$-spaces include the special case of $p=\infty$. Nevertheless, we temporarily exclude this extreme case in order to simplify the comprehension of functions solely in terms of integrations.
[…]
$\|u(t,x)\|_{L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}:=\left(\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}|u(t,x)|^{q}W^{q}(t,x)\mathrm{d}x\right)^{p/q}w^{p}(t)\mathrm{d}t\right)^{1/p}<\infty.$
It is obvious that
$L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)=L_{p,q}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tw^{q}(t)W^{q}(t,x)\mathrm{d}x\right).$
In particular,
$L_{p,p}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{p}(t,x)\mathrm{d}x\right)=L_{p,p}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tw^{p}(t)W^{p}(t,x)\mathrm{d}x\right).$
In this case, our space becomes identical to the $L_{p}$-space with the measure $w^{p}(t)W^{p}(t,x)\mathrm{d}t\mathrm{d}x$, i.e.
$L_{p}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)W^{p}(t,x)\mathrm{d}t\mathrm{d}x\right)$ $=L_{p,p}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{p}(t,x)\mathrm{d}x\right)$
$=L_{p,p}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tw^{p}(t)W^{p}(t,x)\mathrm{d}x\right).$
In addition, observe that
$\|u\|_{L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}$ $=\|u\|_{L_{p}\left((0,T),w^{p}(t)\mathrm{d}t;L_{q}\left(\mathbb{R}^{d},W^{q}(t,x)\mathrm{d}x\right)\right)}$
$=\left(\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}|u(t,x)|^{q}W^{q}(t,x)\mathrm{d}x\right)^{p/q}w^{p}(t)\mathrm{d}t\right)^{1/p}$
$=\left(\int_{0}^{T}\left(\int_{\mathbb{R}^{d}}|w(t)W(t,x)u(t,x)|^{q}\mathrm{d}x\right)^{p/q}\mathrm{d}t\right)^{1/p}$
$=\|wWu\|_{L_{p}\left((0,T);L_{q}\left(\mathbb{R}^{d}\right)\right)}.$
To include the extreme case that $p=\infty$ or $q=\infty$, we define
$\|u(t,x)\|_{L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}:=\|wWu\|_{L_{p}\left((0,T);L_{q}\left(\mathbb{R}^{d}\right)\right)}.$
For $p,q\in[1,\infty)$, we remark that the simpler notations
$L_{p,q}\left((0,T)\times\mathbb{R}^{d},w(t)\mathrm{d}tW(t,x)\mathrm{d}x\right):=L_{p,q}\left((0,T)\times\mathbb{R}^{d},(w^{1/p})^{p}(t)\mathrm{d}t(W^{1/q})^{q}(t,x)\mathrm{d}x\right).$
are used on occasion for the sake of consistency with usual notations of weighted $L_{p,q}$-spaces.
We usually identify $L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)$ and $L_{p}\left((0,T),w^{p}(t)\mathrm{d}t;L_{q}\left(\mathbb{R}^{d},W^{q}(t,x)\mathrm{d}x\right)\right)$ even though the iterated measurability does not imply the joint measurability. We also would like to point out that we would not consider the intricate matters between Borel and Lebesgue measurabilities by opting for the completion of a measure space. Furthermore, we adopt a conventional method to establish the equivalence in the almost everywhere sense across all $L_{p,q}$-spaces without delving into significant details. In particular, we do not make a strict distinction between functions defined throughout $(0,T)\times\mathbb{R}^{d}$ and those defined almost everywhere within $(0,T)\times\mathbb{R}^{d}$. Similarly, this relaxed approach extends to functions defined on $\mathbb{R}^{d}$ as well. With these prerequisites in place, we can now introduce a concept of our weighted local $L_{p,q}$-spaces. It is worth noting that this local characteristic is presented in a slightly different manner with respect to the time variable, which will be further explored in Remark 2.20.
Definition 2.15 (Weighted local $L_{p,q}$-spaces).
Let $p,q\in[1,\infty]$.
  1. (i)
    We define $L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)$ as the class of all measurable functions $u$ on $(0,T)\times\mathbb{R}^{d}$ such that
    $\|u\|_{L_{p,q}\left((0,t_{1})\times B_{R},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}:=\|1_{(0,t_{1})}(t)1_{B_{R}}(x)u(t,x)\|_{L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}<\infty$
    for all $0<t_{1}<T$ and $R\in(0,\infty)$, where $B_{R}:=\{x\in\mathbb{R}^{d}:|x|<R\}$ and $1_{B_{R}}$ denotes the indicator function of $B_{R}$, i.e.
    $1_{B_{R}}(x)=\begin{cases}&1\quad x\in B_{R},\\ &0\quad x\in\mathbb{R}^{d}\setminus B_{R}.\end{cases}$
  2. (ii)
    We define $L_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)$ as the class of all measurable functions $u$ on $(0,T)\times\mathbb{R}^{d}$ such that
    $\|u(t,x)\|_{L_{p,q}\left((0,t_{1})\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)}<\infty$
    for all $0<t_{1}<T$.
We now turn our attention to a significant subset within $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which comprises functions $u$ whose Fourier transform is in an weighted local $L_{p,q}$-space. This subset can be formally expressed as:
$\left\{u:\mathcal{F}[u(t,\cdot)](\xi)\in L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,x)\mathrm{d}x\right)\right\}.$ (2.7)
However, precisely defining the expression above (2.7) poses challenges due to the potential degeneracy of our weight function $W(t,\xi)$ for certain $\xi\in\mathbb{R}^{d}$. It requires employing sophisticated techniques involving identifications with complex-valued functions, as illustrated in (2.3).
Let $u\in L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. Then for each $t\in(0,T)$, we have $\mathcal{F}[u(t)]\in\mathcal{D}^{\prime}(\mathbb{R}^{d})$. Additionally, assume that for each $t\in(0,T)$, $\mathcal{F}[u(t)]$ has a realization on an open set $\mathcal{U}_{t}$ covering the support of $W(t,\cdot)$, i.e. $\overline{\{\xi\in\mathbb{R}^{d}:W(t,\xi)\neq 0\}}\subset\mathcal{U}_{t}$, where $\overline{\{\xi\in\mathbb{R}^{d}:W(t,\xi)\neq 0\}}$ denotes the closure of $\{\xi\in\mathbb{R}^{d}:W(t,\xi)\neq 0\}$. Thus by identifying $\mathcal{F}[u(t)]$ with the realization on $\mathcal{U}$, we may assume that $(t,\xi)\mapsto\mathcal{F}[u(t,\cdot)](\xi):=\mathcal{F}[u](t,\xi):=\mathcal{F}[u(t)](\xi)$ is a complex-valued function defined on $\{(t,\xi)\in(0,T)\times\mathbb{R}^{d}:(t,\xi)\in(0,T)\times\mathcal{U}_{t}\}$. In particular, the term $\|wW\mathcal{F}[u]\|_{L_{p}\left((0,T);L_{q}\left(\mathbb{R}^{d}\right)\right)}$ perfectly makes sense for all $p,q\in[1,\infty]$ by letting the product $wW\mathcal{F}[u]$ be zero outside the set $\{(t,\xi)\in(0,T)\times\mathbb{R}^{d}:(t,\xi)\in(0,T)\times\mathcal{U}_{t}\}$ since $W(t,\xi)=0$ for all $(t,\xi)\notin(0,T)\times\mathcal{U}_{t}$. Therefore simply we may say that
$\mathcal{F}[u(t,\cdot)](\xi)\in L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right),$
if and only if $u$ is an element in $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ so that for each $t\in(0,T)$, $\mathcal{F}[u(t)]$ has a realization on an open set $\mathcal{U}_{t}$ covering the support of $W(t,\cdot)$ and $\|wW\mathcal{F}[u]\|_{L_{p}\left((0,T);L_{q}\left(\mathbb{R}^{d}\right)\right)}<\infty$. Finally, we define subspaces of $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ consisting of $u$ whose Fourier transform lies in a weighted (local) $L_{p,q}$-spaces in the following definition based on the aforementioned criteria.
Definition 2.16.
  1. (i)
    We define subclasses of $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$ consisting of elements whose Fourier transforms with respect to the space variable are in weighted (local) $L_{p,q}$-spaces. We write
    $u\in\mathcal{F}^{-1}L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right),$
    $u\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right),$
    and
    $u\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)$
    if
    $\mathcal{F}[u(t,\cdot)](\xi)\in L_{p,q}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(\xi)\mathrm{d}\xi\right),$
    $\mathcal{F}[u(t,\cdot)](\xi)\in L_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right),$
    and
    $\mathcal{F}[u(t,\cdot)](\xi)\in L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right),$
    respectively. In other words,
    $u\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}x\right)$
    if and only if
    $\|\mathcal{F}[u]\|_{L_{p,q}\left((0,t_{1})\times B_{R},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)}:=\|1_{B_{R}}(\xi)\mathcal{F}[u(t,\cdot)](\xi)\|_{L_{p,q}\left((0,t_{1})\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)}<\infty$
    for all $t_{1}\in(0,T)$ and $R\in(0,\infty)$.
The weights $w^{p}(t)$, $W^{q}(t,x)$, and $W^{q}(t,\xi)$ in the notations are skipped if they are merely $1$. If all weights are $1$, then we also omit $\mathrm{d}t\mathrm{d}x$ and $\mathrm{d}t\mathrm{d}\xi$ in the notations.
These local spaces are employed to give weaker conditions on symbols and find a class of solutions to (1.5) where the existence and uniqueness are guaranteed. Recall Definition 2.9. Then it is obvious that
$L_{1,t\text{-loc}}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)=\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right).$
In addition, by Definition 2.6 and Hölder’s inequality, we have the inclusion that
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\subset\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\subset L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$
for all $p,q\in[1,\infty]$. However,
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)$ $\not\subset\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ (2.8)
in general, since even there is no strict positivity assumption on the weights $w$ and $W$.
Remark 2.17.
There are numerous mathematical conveniences when we exclusively consider positive weights. Nevertheless, it is crucial not to disregard the possibility of weights becoming zero, as this allows us to include the interesting cases in which the symbol $\psi(t,\xi)$ may become degenerate. Furthermore, it is true that, for all values of $p$ and $q$ within the range of $[1,\infty]$,
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)\subset L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$
due to the definition of $\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)$. If both weights $w$ and $W$ are strictly positive, then
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)\subset L_{0}\left((0,T);\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\right)$
[…]
$u\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right).$
However, the space $\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}tW^{q}(t,\xi)\mathrm{d}\xi\right)$ is not a metric space in general if one of the weights is degenerate since $\mathcal{F}[u(t,\cdot)]$ could have any values outside of the supports of the weights as discussed in Remark 2.13.
Remark 2.18.
We skip specifying the exact definitions of corresponding weighted $L_{q}$-spaces treating data given on $\mathbb{R}^{d}$ since they are easily induced from corresponding weighted $L_{p,q}$-spaces on $(0,T)\times\mathbb{R}^{d}$. For instance, $\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d})$ denotes the subspace of $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ so that
$\|\mathcal{F}[u]\|_{L_{q}(B_{R})}<\infty\quad\forall R\in(0,\infty).$
Remark 2.19.
Let $\mathscr{E}^{\prime}(\mathbb{R}^{d})$ represent the dual space of $C^{\infty}(\mathbb{R}^{d})$. Then, for any element $u$ belonging to $\mathscr{E}^{\prime}(\mathbb{R}^{d})$, its Fourier transform, denoted by $\mathcal{F}[u]$, can be interpreted as a function that is locally bounded. To clarify, for every $\xi$ within $\mathbb{R}^{d}$, the function $\mathrm{e}^{-ix\cdot\xi}$ is a smooth function with respect to the variable $x$, and the mapping from $\xi$ to $\langle u,\mathrm{e}^{-ix\cdot\xi}\rangle$ is a continuous function. Moreover, it is widely recognized that this space $\mathscr{E}^{\prime}(\mathbb{R}^{d})$ corresponds to a subset of $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ with compact supports, as discussed in [10, Section 2.3]. As a result, we can conclude that $\mathscr{E}^{\prime}(\mathbb{R}^{d})$ is contained within $\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d})$ for all $q\in[1,\infty]$.
Remark 2.20.
We can describe a function $u$ belonging to the space $L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ as having a strong local integrability with respect to the variable $t$ and a weaker form of local integrability with respect to the variable $x$. To clarify further, consider a topological space denoted by $X$, and recall that when we say a property holds locally, it means that the property is satisfied for all compact subsets $K\subset X$. In this context, our function $u\in L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ can be described as locally integrable with respect to $x$, but it satisfies a somewhat stronger condition compared to the conventional notion of local integrability concerning the time variable $t$. This distinction is why we employed different notations, namely, “$loc$” and ‘$\ell oc$”, to indicate the local integrability properties of $t$ and $x$, respectively.
We finally reach our main theorem.
Theorem 2.21.
Let $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$. Suppose that for all $t\in(0,T)$ and $R\in(0,\infty)$,
$\int_{B_{R}}\exp\left(\int_{0}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi+\int_{B_{R}}\int_{0}^{t}\exp\left(\int_{s}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$ <<$\infty$,< ∞ , (2.9)
where $B_{R}:=\{\xi\in\mathbb{R}^{d}:|\xi|<R\}$. Additionally, assume that for all $t\in(0,T)$ and $R\in(0,\infty)$,
$\int_{B_{R}}\int_{0}^{t}|\psi(\rho,\xi)|\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\rho\mathrm{d}\xi$
$+\int_{B_{R}}\int_{0}^{t}|\psi(\rho,\xi)|\int_{0}^{\rho}\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\rho\mathrm{d}\xi<\infty.$ (2.10)
Then there exists a unique Fourier-space weak solution
$u\in\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$
to (1.5), which is given by
$\langle u(t,\cdot),\varphi\rangle$ $=\int_{\mathbb{R}^{d}}\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathcal{F}[\varphi](\xi)\mathrm{d}\xi$
$\quad+\int_{\mathbb{R}^{d}}\left(\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s\right)\mathcal{F}[\varphi](\xi)\mathrm{d}\xi\quad\forall\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}).$ (2.11)
The proof of Theorem 2.21 is given in Section 5.
Remark 2.22.
The solution $u$ to (1.5) in Theorem 2.21 is not defined as a complex-valued function on $[0,T)\times\mathbb{R}^{d}$. It is a function in the class $L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, which is clarified by the action in (2.11). Therefore, formally, we may say that
$u(t,x)=\mathcal{F}^{-1}\left[\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]\right](x)+\mathcal{F}^{-1}\left[\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s\right](x)$ (2.12)
is a solution to (1.5). It is obvious that this representation for the solution $u$ is identical to the classical one if $\psi$, $u_{0}$, and $f$ satisfy nice mathematical conditions by considering convolutions with kernels as follow :
$u(t,x)=\int_{\mathbb{R}^{d}}K(0,t,x)u_{0}(x-y)\mathrm{d}y+\int_{0}^{t}\int_{\mathbb{R}^{d}}K(s,t,y)f(s,x-y)\mathrm{d}y\mathrm{d}s,$
where the kernel $K(s,t,x)$ is given by
$K(s,t,x)=\mathcal{F}^{-1}\left[\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\right](x).$
This kernel representation shows that it is not hopeful to expect that $u$ has a realization on $[0,T)\times\mathbb{R}^{d}$ if the symbol $\psi(t,\xi)$ is not negative.
Remark 2.23.
We intentionally split the mathematical assumptions in Theorem 2.21 into two parts, (2.9) and (2.10), since the condition (2.10) is sufficient to guarantee the uniqueness of a solution, which is specified in Theorem 4.3. On the other hand, (2.9) guarantee the function $u(t,x)$ defined as (2.12) is well-defined as $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$-valued function. A naive sufficient condition for both (2.9) and (2.10) is a local bounded property on the symbol $\psi(t,\xi)$. More precisely, if $\left\|\psi(\rho,\xi)\right\|_{L_{\infty,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$ for all $t\in(0,T)$ and $R\in(0,\infty)$, then both (2.9) and (2.10) are satisfied, which will be relaxed in Corollary 5.1 with positive weights and general exponents. We also want to mention that both (2.9) and (2.10) are given by combinations of properties on the symbol $\psi$ and data $u_{0}$ and $f$ in a complicated way on purpose to emphasize that the condition on the symbol could be weaken if data $u_{0}$ and $f$ are in better spaces. One of the good examples showing weakening condition on the symbol $\psi$ by enhancing the condition on data $u_{0}$ and $f$ is presented in Corollary 5.3. Additionally, these complicated conditions (2.9) and (2.10) make us handle logarithmic operators instead of merely considering locally bounded symbols, which will be specified in Theorem 2.27 below.
Remark 2.24.
We revisit to mention the importance of the extension of the inverse Fourier transform to all distributions. Assume that the symbol $\psi(t,\xi)$ is locally bounded. Then the function $\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]$ is merely locally integrable with respect to $\xi$. Thus, there is no assurance that $\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]$ qualifies as a tempered distribution on $\mathbb{R}^{d}$. Consequently, the function $u$ in equation (2.12) cannot be defined because the inverse Fourier transform is typically applicable only to tempered distributions on $\mathbb{R}^{d}$ (if we do not consider an extension to $\mathbb{C}^{d}$). However, it is important to note that we extended the inverse Fourier transform to all distributions as an element in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$. Therefore, the function $u$ is well-defined as a new extension of the inverse Fourier transform to a distribution even if the functions $\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]$ and $\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s$ are merely locally integrable. Especially, it is easy to check that
$u\in\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$
due to (2.9) and (2.10).
Remark 2.25.
It is not trivial to understand realizations of data $u_{0}$ and $f$ since all data $u_{0}$ and $f$ are given based on Fourier transforms acting on $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$. However, it is important to note that $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ is a broader class than $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Within $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, there exist various well-known subspaces that help characterize regularities of elements in $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. One notable example is the Bessel potential space. Given that $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ is larger than any subspace of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, we could apply our theorem to data $u_{0}$ and $f$ residing in Bessel potential spaces. We suggest a weighted Bessel potential space as an important example of realizations of data $u_{0}$ and $f$ in the next section. The specific results will be explicitly presented in Corollaries 5.4, 5.6, and 5.7.
It is readily verifiable that the logarithmic Laplacian operator can be regarded as a specific example of the operator $\psi(t,-i\nabla)$. Recall that the logarithmic Laplacian operator could be defined as follows
$\mathcal{F}[\log(-\Delta)u]=\log|\xi|^{2}\mathcal{F}[u]$
for a nice function $u$. This means that the logarithmic Laplacian operator $\log(-\Delta)$ is a pseudo-differential operator characterized by the symbol $\log|\xi|^{2}$. The logarithmic Laplacian operator is an intriguing example of $\psi(t,-i\nabla)$ because its symbol is negative for $|\xi|<1$ and positive for $|\xi|>1$. Additionally, the symbol is not a polynomial. In other words, the logarithmic Laplacian operator is a simple illustration of a pseudo-differential operator with a non-polynomial symbol that changes its sign. As an application of Theorem 2.21, we explore more generalized logarithmic-type operators with complex-valued coefficients, which naturally encompass the logarithmic Laplacian operator. We now present an evolution equation involving a logarithmic-type operator as a specific case of (1.5) below:
$\begin{cases}\partial_{t}u(t,x)=\beta(t)\log\left(\psi_{exp}(t,-i\nabla)\right)u(t,x)+f(t,x),\quad&(t,x)\in(0,T)\times\mathbb{R}^{d},\\ u(0,x)=u_{0},\quad&x\in\mathbb{R}^{d},\end{cases}$ (2.13)
where
$\log\left(\psi_{exp}(t,-i\nabla)\right)u(t,x):=\mathcal{F}^{-1}\left[\log\left(\psi_{exp}(t,\xi)\right)\mathcal{F}[u(t,\cdot)]\right](x),$ (2.14)
$\beta(t)$ is a complex-valued measurable function defined on $(0,T)$, and $\psi_{exp}(t,\xi)$ is a complex-valued measurable function defined on $(0,T)\times\mathbb{R}^{d}$. In particular, if $\beta(t)=1$ and $\psi_{exp}(t,\xi)=|\xi|^{2}$ for all $t$, then
$\beta(t)\log\left(\psi_{exp}(t,-i\nabla)\right)u(t,x)=\log(-\Delta)u(t,x).$
A comprehensive understanding of a weak solution $u$ to the equation (2.13) can be achieved by referring to Definition 2.10. Moreover, we assume
$\psi_{exp}(t,\xi)\neq 0\quad(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d}.$ (2.15)
It is essential for the right-hand side of (2.14) to be meaningful that the condition stated in (2.15) is satisfied. This requirement is necessary due to the presence of the logarithm’s domain.
Remark 2.26.
Let’s define $\psi(t,\xi)$ as the natural logarithm of $\psi_{exp}(t,\xi)$, i.e.
$\psi(t,\xi)=\log\left(\psi_{exp}(t,\xi)\right).$
From a heuristic standpoint, we can express this relationship as
$\exp\left(\psi(t,\xi)\right)=\psi_{exp}(t,\xi),$
which explains the reason using the notation $\psi_{exp}(t,\xi)$. In other words, the subscript in $\psi_{exp}(t,\xi)$ signifies the exponential operation.
We restate a theorem opting for (2.13), which is one of the major applications of Theorem 2.21.
Theorem 2.27.
Let $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$. Assume that (2.15) holds and
$\int_{0}^{t}|\beta(s)|\mathrm{d}s<\infty\quad\forall t\in(0,T).$ (2.16)
Additionally, suppose that
$\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\frac{1}{t-s}\int_{s}^{t}|\psi_{exp}(r,\xi)|^{(t-s)\Re[\beta(r)]}\mathrm{d}r\right)<\infty$ (2.17)
and
$\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\int_{s}^{t}|\beta(\rho)|\left(1+\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|\right)\left(\frac{1}{\rho-s}\int_{s}^{\rho}\left(|\psi_{exp}(r,\xi)|^{(\rho-s)\Re[\beta(r)]}\right)\mathrm{d}r\right)\mathrm{d}\rho\right)<\infty$ (2.18)
for all $t\in(0,T)$ and $R\in(0,\infty)$. Then there exists a unique Fourier-space weak solution $u$ to (2.13) in
$\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},|\beta(t)|\mathrm{d}t|\log\left(\psi_{exp}(t,\xi)\right)|\mathrm{d}\xi\right).$
The proof of this theorem will be given in Section 6.
Remark 2.28.
It is notable that the complex-valued function $\psi_{exp}(t,\xi)$ could have a negative real value. It is possible because we have not imposed any regularity conditions on the symbol $\log\left(\psi_{exp}(t,\xi)\right)$, which means there is no need to consider the analytic continuation of the natural logarithm.
Remark 2.29.
The condition (2.18) could be replaced by
$\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\int_{s}^{t}|\beta(\rho)|\left|\log\left(\psi_{\exp}(\rho,\xi)\right)\right|\left(\frac{1}{\rho-s}\int_{s}^{\rho}\left(|\psi_{exp}(r,\xi)|^{(\rho-s)\Re[\beta(r)]}\right)\mathrm{d}r\right)\mathrm{d}\rho\right)<\infty$
since it is obvious that
$\left(1+\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|\right)\approx\left|\log\left(\psi_{\exp}(\rho,\xi)\right)\right|$
by considering a branch cut.
Remark 2.30.
Consider the logarithmic Laplacian operator
$\log(-\Delta)u(t,x)$
as a special case of
$\beta(t)\log\left(\psi_{exp}(t,-i\nabla)\right)u(t,x)$
in Theorem 2.27. It is easy to confirm that (2.17) is valid for this simple case since $\beta(t)=1$ and $\psi_{exp}(t,\xi)=|\xi|^{2}$ for all values of $t$ and $\xi$. Nonetheless, it might seem that (2.18) is not straightforward. However, it can be readily established through elementary calculations, as demonstrated in the proof of Corollary 6.1. Furthermore, we present a straightforward extension of this result for dealing with operators that result from a composition involving the $\log$ function and a general second-order differential operator, which is detailed in Corollary 6.1.
Next, we consider another important case handling a second-order differential operator without ellipticity. We investigate a second-order evolution equation with complex-valued coefficients $a^{ij}(t)$, $b^{i}(t)$, and $c(t)$ as follows:
$\begin{cases}\partial_{t}u(t,x)=a^{ij}(t)u_{x^{i}x^{j}}(t,x)+b^{j}(t)u_{x^{j}}(t,x)+c(t)u(t,x)+f(t,x),\quad&(t,x)\in(0,T)\times\mathbb{R}^{d},\\ u(0,x)=u_{0},\end{cases}$ (2.19)
where Einstein’s summation convention is being enforced here. Suppose that all coefficients $a^{ij}(t)$, $b^{j}(t)$, and $c(t)$ are measurable and defined on $(0,T)$. Formally, due to some properties of the Fourier transform and the inverse Fourier transform,
$a^{ij}(t)u_{x^{i}x^{j}}(t,x)+b^{j}(t)u_{x^{j}}(t,x)+c(t)u(t,x)=\mathcal{F}^{-1}\left[\left(-a^{ij}(t)\xi^{i}\xi^{j}+ib^{j}(t)+c(t)\right)\mathcal{F}[u(t,\cdot)](\xi)\right](x).$
Hence, equation (2.19) can be seen as a specific instance of equation (1.5) when considering the symbol:
$\psi(t,\xi)=-a^{ij}(t)\xi^{i}\xi^{j}+ib^{j}(t)\xi^{j}+c(t),$ (2.20)
where $i$ beside $b^{j}(t)$ is the imaginary number so that $i^{2}=-1$. To maintain clarity and facilitate readers’ comprehension, we provide a modified version of the definition of a Fourier-space weak solution for second-order cases. This indicates that the structure of our weak formulation closely parallels that of the classical one.
Definition 2.31 (A weak solution tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$).
Let $u\in L_{0}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, $u_{0}\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, and $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. Then we say that $u$ is a weak solution (tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$) to (2.19) if for any $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$,
$\langle u(t,\cdot),\varphi\rangle=\langle u_{0},\varphi\rangle+\int_{0}^{t}\left(a^{ij}(s)\langle u(s,\cdot),\varphi_{x^{i}x^{j}}\rangle-b^{j}(s)\langle u(s,\cdot),\varphi_{x^{j}}\rangle+c(s)\langle u(s,\cdot),\varphi\rangle\right)\mathrm{d}s+\int_{0}^{t}\langle f(s,\cdot),\varphi\rangle\mathrm{d}s$
$(a.e.)~{}t\in(0,T).$ (2.21)
Remark 2.32.
Let us reiterate that a test function is conventionally given by a smooth function with compact support in a weak formulation. In simpler terms, we commonly say that $u$ qualifies as a (classical) weak solution to equation (2.19) if (2.21) is valid for all $\varphi\in\mathcal{D}(\mathbb{R}^{d})$ instead of $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$. To ensure the meaningfulness of this traditional weak formulation, it is essential for $u$ to be valued in $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ and defined almost everywhere on $(0,T)$. Sufficiently assume that $u$ is a $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued measurable function on $(0,T)$ and it is a classical weak solution. Then this classical solution $u$ becomes identical with the weak solution (tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$) in Definition 2.31 since $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$ is dense within $\mathcal{S}(\mathbb{R}^{d})$ concerning the topology defined by the Schwartz semi-norms. However, it is crucial to highlight that for a solution $u$ to equation (2.19), $u(t,\cdot)$ does not belong to $\mathcal{D}^{\prime}(\mathbb{R}^{d})$ (and thus $u(t,\cdot)\notin\mathcal{S}^{\prime}(\mathbb{R}^{d})$) if there is no ellipticity in the leading coefficients $a^{ij}(t)$. This emphasizes the necessity for a novel concept of a weak solution, such as the Fourier-space weak solution.
Remark 2.33.
The weak solution (tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$) is equivalent to the Fourier-space weak solution introduced in Definition 2.10 for the particular symbol $\psi$ defined in (2.20) if the Fourier transform of $u$ is locally integrable (with respect to the space variable) since
$a^{ij}(s)\langle u(s,\cdot),\varphi_{x^{i}x^{j}}\rangle-b^{j}(s)\langle u(s,\cdot),\varphi_{x^{j}}\rangle+c(s)\langle u(s,\cdot),\varphi\rangle$
$=a^{ij}(s)\left(\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi_{x^{i}x^{j}}]\right)_{L_{2}(\mathbb{R}^{d})}-b^{j}(s)\left(\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi_{x^{j}}]\right)_{L_{2}(\mathbb{R}^{d})}+c(s)\left(\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}$
$=\left(\left(-a^{ij}(s)\xi^{i}\xi^{j}+ib^{j}(s)\xi^{j}+c(s)\right)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}$
$=\left\langle\psi(s,-i\nabla)u(s,\cdot),\varphi\right\rangle,$
which are easily derived from some properties of the Fourier transform.
We are now prepared to present the theorem regarding equation (2.19). It is important to note that all the conditions specified for the coefficients in the theorem merely involve local integrability properties. Especially, all coefficients could be complex-valued, and the matrix consisting of the leading coefficients $a^{ij}(t)$ does not need to be positive semi-definite.
Theorem 2.34.
Let $p,q\in[1,\infty]$, $W_{0}(\xi)$ and $W_{1}(\xi)$ be positive measurable functions on $\mathbb{R}^{d}$, $W_{2}(t,\xi)$ be a positive measurable function on $(0,T)\times\mathbb{R}^{d}$, $u_{0}\in\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d},W_{0}^{q}(\xi)\mathrm{d}\xi)$, and
$f\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tW_{2}^{q}(t,\xi)\mathrm{d}\xi\right).$
Assume that $a^{ij},b^{j},c\in L_{p,loc}\left((0,T)\right)$ for all $i,j\in\{1,\ldots,d\}$, i.e.
$\sum_{i=1}^{d}\sum_{j=1}^{d}\left(\|a^{ij}\|_{L_{p}\left((0,t)\right)}+\|b^{j}\|_{L_{p}\left((0,t)\right)}+\|c\|_{L_{p}\left((0,t)\right)}\right)<\infty\quad\forall t\in(0,T).$ (2.22)
Additionally, suppose that $W_{1}$ is a lower bound of $W_{2}$ and both $W_{0}$ and $W_{1}$ have local lower bounds, i.e. for each $t\in(0,T)$ and $R\in(0,\infty)$, there exist positive constants $\kappa_{0}(R)$ and $\kappa_{1}(R)$ so that
$W_{0}(\xi)\geq\kappa_{0}(R)\quad\forall\xi\in B_{R}$ (2.23)
and
$W_{2}(s,\xi)\geq W_{1}(\xi)\geq\kappa_{1}(R)\quad\forall(s,\xi)\in(0,t)\times B_{R}.$ (2.24)
Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t\left(1+|a^{ij}(t)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)|\right)\mathrm{d}\xi\right)$
and the solution $u$ is given by
$u(t,x)$ $=\mathcal{F}^{-1}\left[\exp\left(\int_{0}^{t}\left(-a^{ij}(r)\xi^{i}\xi^{j}+ib^{j}(r)\xi^{j}+c(r)\right)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\right](x)$
$\quad+\mathcal{F}^{-1}\left[\int_{0}^{t}\exp\left(\int_{s}^{t}\left(-a^{ij}(r)\xi^{i}\xi^{j}+ib^{j}(r)\xi^{j}+c(r)\right)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s\right](x).$
Here the solution $u$ is given formally as discussed in Remark 2.22.
We will provide the proof of this theorem in Section 7. It is worth noting that when dealing with data $u_{0}$ and $f$ in weighted Bessel potentials, this theorem can be readily applied. We demonstrate these applications as corollaries of Theorem 2.34 in Section 7.
3. Fourier transforms and weighted Bessel potential spaces
In this section, our objective is to present two types of weighted Bessel potential spaces. These spaces exhibit numerous intriguing properties that can be effectively characterized through the use of Fourier and inverse Fourier transforms. To begin, we recall the definitions and some properties of the Fourier and inverse Fourier transforms. For a measurable function $f$ on $\mathbb{R}^{d}$, we denote the $d$-dimensional Fourier transform of $f$ by
$\mathcal{F}[f](\xi):=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{-i\xi\cdot x}f(x)\mathrm{d}x$
and the $d$-dimensional inverse Fourier transform of $f$ by
$\mathcal{F}^{-1}[f](x):=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}f(\xi)\mathrm{d}\xi.$
Due to the Parseval-Plancherel identity,
$\|f\|_{L_{2}(\mathbb{R}^{d})}=\|\mathcal{F}[f]\|_{L_{2}(\mathbb{R}^{d})}=\|\mathcal{F}^{-1}[f]\|_{L_{2}(\mathbb{R}^{d})}\quad\forall f\in L_{1}(\mathbb{R}^{d})\cap L_{2}(\mathbb{R}^{d}).$
The above equalities imply that the Fourier and inverse Fourier transforms are $L_{2}(\mathbb{R}^{d})$-isometries by considering $L_{2}$-extensions based on completeness of $L_{2}(\mathbb{R}^{d})$. Additionally, due to the Riesz–Thorin theorem, there exist the extensions of the Fourier and inverse Fourier transforms which are bounded from $L_{1}(\mathbb{R}^{d})+L_{2}(\mathbb{R}^{d})$ to $L_{\infty}(\mathbb{R}^{d})+L_{2}(\mathbb{R}^{d})$. In particular, for any $\theta\in[0,1]$,
$\|\mathcal{F}[f]\|_{L_{2/\theta}}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(1-\theta)}{2}}\|f\|_{L_{2/(2-\theta)}}\quad\forall f\in L_{2/(2-\theta)}$
and
$\|\mathcal{F}^{-1}[f]\|_{L_{2/\theta}}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(1-\theta)}{2}}\|f\|_{L_{2/(2-\theta)}}\quad\forall f\in L_{2/(2-\theta)},$
where $\frac{2}{0}:=\infty$. Therefore, for any $p\in[1,2]$ and $f\in L_{p}(\mathbb{R}^{d})$, it is possible to regard the Fourier transform and the inverse Fourier transform of $f$ as functions within $L_{p^{\prime}}(\mathbb{R}^{d})$, where $p^{\prime}=\frac{p}{p-1}$ and $\frac{1}{0}:=\infty$. This is a crucial foundation for our theory, as it allows us to address our solutions and data in a strong sense. To be more specific, for any $p\in[2,\infty]$ and any function $f$ in $L_{p^{\prime}}(\mathbb{R}^{d})$, we have the following inequalities:
$\|\mathcal{F}[f]\|_{L_{p}}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(p-2)}{2p}}\|f\|_{L_{p^{\prime}}}$ (3.1)
and
$\|\mathcal{F}^{-1}[f]\|_{L_{p}}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(p-2)}{2p}}\|f\|_{p^{\prime}}.$ (3.2)
We now turn our attention to the Fourier and inverse Fourier transforms applied to the space of tempered distributions on $\mathbb{R}^{d}$. These transformations are established in a weak sense, making use of the advantageous properties of the Schwartz class $\mathcal{S}(\mathbb{R}^{d})$. Suppose we have a function $f$ belonging to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. In that case, we define the Fourier and inverse Fourier transforms as follows:
$\langle\mathcal{F}[f],\varphi\rangle=\langle f,\mathcal{F}^{-1}[\varphi]\rangle\quad\forall\varphi\in\mathcal{S}(\mathbb{R}^{d})$
and
$\langle\mathcal{F}^{-1}[f],\varphi\rangle=\langle f,\mathcal{F}[\varphi]\rangle\quad\forall\varphi\in\mathcal{S}(\mathbb{R}^{d}).$
Here, the notation $\langle f,\mathcal{F}^{-1}[\varphi]\rangle$ represents the duality pairing, which means that $\langle f,\mathcal{F}^{-1}[\varphi]\rangle=f\left(\mathcal{F}^{-1}[\varphi]\right)$ is the value assigned to $\mathcal{F}^{-1}[\varphi]$ under the action of the linear functional $f$. It is important to note that the Fourier and inverse Fourier transforms of $f$ once again become tempered distributions on $\mathbb{R}^{d}$, thanks to the well-established properties of functions within the Schwartz class. Furthermore, it is worth noting that $(1+|\cdot|^{2})^{\gamma_{2}/2}f$ remains a tempered distribution as well if we define the action on test functions as follows:
$\left\langle(1+|\cdot|^{2})^{\gamma_{2}/2}f,\varphi\right\rangle:=\left\langle f,(1+|\cdot|^{2})^{\gamma_{2}/2}\varphi(\cdot)\right\rangle\quad\text{for all}\quad\varphi\in\mathcal{S}(\mathbb{R}^{d}).$
This is possible because the function $(1+|x|^{2})^{\gamma_{2}/2}\varphi(x)$ belongs to the Schwartz class $\mathcal{S}(\mathbb{R}^{d})$ for any $\varphi\in\mathcal{S}(\mathbb{R}^{d})$. As a result, these operations on tempered distributions enable us to establish the Bessel potential spaces.
Let $\gamma\in\mathbb{R}$ and $p\in[1,\infty]$. Recall that we define the Bessel potential space with order $\gamma$ and exponent $p$, denoted by $H_{p}^{\gamma}(\mathbb{R}^{d})$, as the class of all tempered distributions $f$ on $\mathbb{R}^{d}$ satisfying the condition:
$(I-\Delta)^{\gamma/2}f:=\mathcal{F}^{-1}\left[\left(1+|\cdot|^{2}\right)\mathcal{F}[f]\right]\in L_{p}(\mathbb{R}^{d}).$
For a more comprehensive understanding, further details on these concepts can be found in references such as [10] and [8]. Now, we proceed to introduce two generalizations of the Bessel potential spaces with additional weighting factors below. We call these spaces an inner weighted Bessel potential space and an outer weighted Bessel potential space, respectively. It is worth noting that we have not found any previous literature that discusses these types of weighted Bessel potential space in general. However, certain outer weighted Bessel potential spaces can be regarded as particular cases of weighted Bessel potential spaces with Muckenhoupt’s weight. We will provide further clarification of this relationship before presenting Proposition 3.5 later on. We are now ready to present the definitions for both inner and outer weighted Bessel potential spaces.
Definition 3.1 (Inner weighted Bessel potential space).
Let $\gamma_{1},\gamma_{2}\in\mathbb{R}$ and $p\in[1,\infty]$. We use the notation $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ to denote the space of all tempered distributions $f$ on $\mathbb{R}^{d}$ such that
$(I-\Delta)^{\gamma_{1}/2}\left((1+|\cdot|^{2})^{\gamma_{2}/2}f\right):=\mathcal{F}^{-1}\left[\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\mathcal{F}\left[(1+|\cdot|^{2})^{\gamma_{2}/2}f\right]\right]\in L_{p}(\mathbb{R}^{d}),$
which implies that there exists a $g\in L_{p}(\mathbb{R}^{d})$ such that
$\left\langle\mathcal{F}^{-1}\left[\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\mathcal{F}\left[\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f\right]\right],\varphi\right\rangle$ $:=\left\langle f,\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}\varphi\right\rangle$
$=\int_{\mathbb{R}^{d}}g(x)\varphi(x)\mathrm{d}x\qquad\forall\varphi\in\mathcal{S}(\mathbb{R}^{d}).$
Then $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ becomes a Banach space with the norm
$\|f\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}:=\left\|(I-\Delta)^{\gamma_{1}/2}\left((1+|\cdot|^{2})^{\gamma_{2}/2}f\right)\right\|_{L_{p}(\mathbb{R}^{d})}.$
We say that $\gamma_{1}$, $\gamma_{2}$, and $p$ are regularity exponent, weight exponent, and integrability exponent of $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
Definition 3.2 (Outer weighted Bessel potential space).
Let $\gamma_{1},\gamma_{2}\in\mathbb{R}$ and $p\in[1,\infty]$. We use the notation $H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ to denote the space of all tempered distributions $f$ on $\mathbb{R}^{d}$ such that
$(1+|\cdot|^{2})^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f\in L_{p}(\mathbb{R}^{d}).$
Then $H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ becomes a Banach space with the norm
$\|f\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}:=\left\|\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f\right\|_{L_{p}(\mathbb{R}^{d})}.$
Same terminologies are used for exponents $\gamma_{1}$, $\gamma_{2}$, and $p$. In other words, $\gamma_{1}$, $\gamma_{2}$, and $p$ are called regularity exponent, weight exponent, and integrability exponent of $H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
We would not delve into the specific details demonstrating that these two spaces are indeed Banach spaces. This can be readily proved by relying on the completeness of $L_{p}$-spaces. In addition, it is obvious that
$H_{p}^{\gamma_{1}}(\mathbb{R}^{d})=H_{p,in}^{\gamma_{1},0}(\mathbb{R}^{d})=H_{p,out}^{\gamma_{1},0}(\mathbb{R}^{d})$
and
$H_{p,in}^{0,\gamma_{2}}(\mathbb{R}^{d})=H_{p,out}^{0,\gamma_{2}}(\mathbb{R}^{d}).$
Remark 3.3.
Let $\gamma_{1}>0$. It is a well-known fact that the Bessel potentials $(I-\Delta)^{-\gamma_{1}/2}$ have elegant representations in terms of Green functions (as discussed in [9, Section 1.2.2]). In essence, this means that there exists a smooth function $G_{\gamma_{1}}(x)$ defined on $\mathbb{R}^{d}$ except for the origin, such that
$(I-\Delta)^{-\gamma_{1}/2}g(x)=\int_{\mathbb{R}^{d}}G_{\gamma_{1}}(y)g(x-y)\mathrm{d}y.$
for any function $g$ in $L_{p}(\mathbb{R}^{d})$. For this particular case, it is possible to understand realizations of all tempered distributions within these weighted spaces solely through integrals. In other words, a tempered distribution $f$ on $\mathbb{R}^{d}$ belongs to $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ if and only if there exists a function $g$ in $L_{p}(\mathbb{R}^{d})$ such that
$f(x)=(1+|x|^{2})^{-\gamma_{2}/2}\int_{\mathbb{R}^{d}}G_{\gamma_{1}}(y)g(x-y)\mathrm{d}y.$
Similarly, $f\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ if and only if there exist a $g\in L_{p}(\mathbb{R}^{d})$ such that
$f(x)=\int_{\mathbb{R}^{d}}G_{\gamma_{1}}(y)(1+|x-y|^{2})^{-\gamma_{2}/2}g(x-y)\mathrm{d}y.$
The function $G_{\gamma_{1}}(x)$ is given by
$G_{\gamma_{1}}(x)=\frac{1}{\Gamma(\frac{\gamma_{1}}{2})}\int_{0}^{\infty}\mathrm{e}^{-t}\mathrm{e}^{-|x|^{2}/(4t)}t^{(\gamma_{1}-d)/2}\frac{\mathrm{d}t}{t},$
where $\Gamma$ denotes the Gamma function. Additionally, the behaviors of $G$ is characterized as follows ([9, Proposition 1.2.5]): For large values of $|x|$, it satisfies
$G_{\gamma_{1}}(x)\lesssim\mathrm{e}^{-|x|^{2}/2}$
and for small values of $|x|$, the behavior of $G_{\gamma_{1}}(x)$ can be described by the following cases according to the dimension $d$:
$G_{\gamma_{1}}(x)\approx\begin{cases}&|x|^{\gamma_{1}-d}+1+O(|x|^{\gamma_{1}-d+2})\quad\text{if $0<\gamma_{1}<d$}\\ &\log\frac{2}{|x|}+1+O(|x|^{2})\quad\text{if $\gamma_{1}=d$}\\ &1+O(|x|^{\gamma_{1}-d})\quad\text{if $\gamma_{1}>d$}.\end{cases}$
We establish a few straightforward embedding inequalities for these two weighted spaces, beginning with the inner spaces. The first embedding presented below is effective concerning the regularity exponents of inner spaces.
Proposition 3.4.
Let $\gamma_{1},\gamma_{2},\tilde{\gamma}_{1}\in\mathbb{R}$, $p\in(1,\infty]$, and $f\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$. Assume that $\gamma_{1}\geq\tilde{\gamma}_{1}$. Then
$\left\|f\right\|_{H_{p,in}^{\tilde{\gamma}_{1},\gamma_{2}}(\mathbb{R}^{d})}\leq\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
Proof.
Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$. Then by Hölder’s inequality and an $L_{p}$-boundedness of the Bessel potential (cf. [9, Section 1.2.2]),
$\left|\left\langle f,\varphi\right\rangle\right|$ $=\left|\left\langle(I-\Delta)^{\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f\right),(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\varphi\right)\right\rangle\right|$
$\leq\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}\left\|(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\varphi\right)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}$
$=\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}\left\|(I-\Delta)^{(\tilde{\gamma}_{1}-\gamma_{1})/2}\left(I-\Delta\right)^{-\tilde{\gamma}_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\varphi\right)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}$
$\leq\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}\left\|\left(I-\Delta\right)^{-\tilde{\gamma}_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\varphi\right)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}.$
Thus considering
$\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}\left(I-\Delta\right)^{\tilde{\gamma}_{1}/2}\varphi$
instead of $\varphi$, we have
$\left|\left\langle\left(I-\Delta\right)^{\tilde{\gamma}_{1}/2}\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f,\varphi\right\rangle\right|$ $=\left|\left\langle f,\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}\left(I-\Delta\right)^{\tilde{\gamma}_{1}/2}\varphi\right\rangle\right|$
$\leq\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}\left\|\varphi\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}.$
Taking the supremum over all $\varphi$ such that $\|\varphi\|_{L_{p^{\prime}}(\mathbb{R}^{d})}=1$, we have
$\left\|f\right\|_{H_{p,in}^{\tilde{\gamma}_{1},\gamma_{2}}(\mathbb{R}^{d})}\leq\left\|f\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
An analogous embedding inequality is valid for outer spaces. This embedding is provided in relation to the weight exponents of outer spaces.
Proposition 3.5.
Let $\gamma_{1},\gamma_{2},\tilde{\gamma}_{2}\in\mathbb{R}$, $p\in(1,\infty]$, and $f\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$. Assume that $\gamma_{2}\geq\tilde{\gamma}_{2}$. Then
$\left\|f\right\|_{H_{p,out}^{\gamma_{1},\tilde{\gamma}_{2}}(\mathbb{R}^{d})}\leq\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
Proof.
Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$. Then
$\left|\left\langle f,\varphi\right\rangle\right|$ $=\left|\left\langle\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f,\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\varphi\right\rangle\right|$
$\leq\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}}\left\|\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\varphi\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}$
$\leq\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}}\left\|\left(1+|\cdot|^{2}\right)^{(\tilde{\gamma}_{2}-\gamma_{2})/2}\left(1+|\cdot|^{2}\right)^{-\tilde{\gamma}_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\varphi\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}$
$\leq\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}}\left\|\left(1+|\cdot|^{2}\right)^{-\tilde{\gamma}_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\varphi\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}.$
Thus
$\left|\left\langle\left(1+|\cdot|^{2}\right)^{\tilde{\gamma}_{2}/2}(I-\Delta)^{\gamma_{1}/2}f,\varphi\right\rangle\right|$ $=\left|\left\langle f,(I-\Delta)^{\gamma_{1}/2}\left(1+|\cdot|^{2}\right)^{\tilde{\gamma}_{2}/2}\varphi\right\rangle\right|$
$\leq\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}}\left\|\varphi\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}.$
Finally, taking the supremum over all $\varphi$ such that $\|\varphi\|_{L_{p^{\prime}}(\mathbb{R}^{d})}=1$, we obtain the result. The proposition is proved. ∎
Remark 3.6.
The case $p=1$ is excluded in both Propositions 3.4 and 3.5 since the dual space of $L_{\infty}(\mathbb{R}^{d})$ is strictly larger than $L_{1}(\mathbb{R}^{d})$.
The two embedding inequalities mentioned above appear quite straightforward, as evidenced by their proofs. However, we will now delve into a non-trivial embedding that relies on a weighted multiplier. To begin, let’s revisit the definition of Muckenhoupt’s weight. We say that a non-negative measurable function $w$ on $\mathbb{R}^{d}$ belongs to $A_{p}$ with $p\in(1,\infty)$ if
$\sup\left(\frac{1}{|B|}\int_{B}w(x)\mathrm{d}x\right)\left(\frac{1}{|B|}\int_{B}w(x)^{-1/(p-1)}\mathrm{d}x\right)^{p-1}<\infty,$
where the supremum is taken over all Euclidean balls on $\mathbb{R}^{d}$ and $|B|$ denotes the Lebesgue measure of $B$. It is well-known that the mapping $\xi\in\mathbb{R}^{d}\mapsto|\xi|^{\gamma}$ is in $A_{p}$ if and only if $\gamma\in(-d,d(p-1))$ (cf. [8, Example 7.1.7]). Thus it is easy to verify that $(1+|\xi|^{2})^{\gamma/2}$ is in $A_{p}$ for all $\gamma\in(-d,d(p-1))$.
It is noteworthy that there is a scarcity of research papers dedicated to the exploration of weighted Bessel potential spaces with Muckenhoupt’s weights, despite these spaces representing natural extensions of weighted Sobolev spaces. Furthermore, handling these spaces becomes relatively straightforward by applying weighted multiplier theories in conjunction with classical Bessel potential spaces.
For additional insights and properties of weighted Bessel potential spaces with Muckenhoupt’s weights, we recommend referring to [1, Appendix] and [23, Section 3]. The proposition we are about to present can be derived as a specific case of the results cited above. Nonetheless, we provide a simple proof to illustrate that it is an easily applicable result within the context of a weighted multiplier theory.
Proposition 3.7.
Let $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $p\in(1,\infty)$, $\gamma_{2}\in\left(-\frac{d}{p},\frac{d(p-1)}{p}\right)$, and $f\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$. Assume that $\gamma_{1}\geq\tilde{\gamma}_{1}$. Then
$\left\|f\right\|_{H_{p,out}^{\tilde{\gamma}_{1},\gamma_{2}}(\mathbb{R}^{d})}\lesssim\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
Proof.
It is obvious that $(1+|\xi|^{2})^{(p\gamma_{2})/2}$ is in $A_{p}$ since $-d<p\gamma_{2}<d(p-1)$. Moreover, it is also easy to show that $(1+|\xi|^{2})^{(\tilde{\gamma}_{1}-\gamma_{1})/2}$ is a weighted $L_{p}$-multiplier since $\gamma_{1}\geq\tilde{\gamma}_{1}$ (cf. [28]). Therefore, we have
$\left\|f\right\|_{H_{p,out}^{\tilde{\gamma}_{1},\gamma_{2}}(\mathbb{R}^{d})}$ $=\left\|\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\tilde{\gamma}_{1}/2}f\right\|_{L_{p}(\mathbb{R}^{d})}$
$=\left\|\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{(\tilde{\gamma}_{1}-\gamma_{1})/2}(I-\Delta)^{\gamma_{1}/2}f\right\|_{L_{p}(\mathbb{R}^{d})}$
$=\left\|(I-\Delta)^{(\tilde{\gamma}_{1}-\gamma_{1})/2}(I-\Delta)^{\gamma_{1}/2}f(x)\right\|_{L_{p}(\mathbb{R}^{d},\left(1+|x|^{2}\right)^{(p\gamma_{2})/2}\mathrm{d}x)}$
$\lesssim\left\|(I-\Delta)^{\gamma_{1}/2}f(x)\right\|_{L_{p}(\mathbb{R}^{d},\left(1+|x|^{2}\right)^{(p\gamma_{2})/2}\mathrm{d}x)}$
$=\left\|\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f\right\|_{L_{p}(\mathbb{R}^{d})}$
$=\left\|f\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
The proposition is proved. ∎
Subsequently, we highlight an evident observation that both Fourier and inverse Fourier transforms operate similarly within these weighted spaces. This observation will be employed extensively in the paper without repeated explicit mention.
Proposition 3.8.
Let $\gamma_{1},\gamma_{2}\in\mathbb{R}$, $p\in[1,\infty]$, and $f$ be a tempered distribution on $\mathbb{R}^{d}$.
  1. (i)
    $\mathcal{F}[f]\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{if and only if}~{}\mathcal{F}^{-1}[f]\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d}).$
    Additionally,
    $\left\|\mathcal{F}[f]\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}=\left\|\mathcal{F}^{-1}[f]\right\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
  2. (ii)
    $\mathcal{F}[f]\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{if and only if}~{}\mathcal{F}^{-1}[f]\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d}).$
    Additionally,
    $\left\|\mathcal{F}[f]\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}=\left\|\mathcal{F}^{-1}[f]\right\|_{H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
Proof.
It is trivial since
$\mathcal{F}[\varphi](\xi)=\mathcal{F}^{-1}[\varphi](-\xi).$
for all $\varphi\in\mathcal{S}(\mathbb{R}^{d})$ and $\xi\in\mathbb{R}^{d}$. ∎
We also investigate relations between inner and outer spaces using Riesz-Thorin inequalities (3.1) and (3.2). The Fourier transform operates as a bridge by interchanging regularity and weight exponents when we connect inner and outer spaces. It is important to note that the range of $p$ naturally becomes restrictive in accordance with the constraints imposed by Riesz-Thorin’s theorem.
Proposition 3.9.
Let $p\in[1,2]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, and $f$ be a tempered distribution on $\mathbb{R}^{d}$.
  1. (i)
    If $f\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, then
    $\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{p^{\prime},out}(\mathbb{R}^{d})}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-p)}{2p}}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,in}(\mathbb{R}^{d})}.$
  2. (ii)
    If $\mathcal{F}[f]\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, then
    $\|f\|_{H^{\gamma_{2},\gamma_{1}}_{p^{\prime},out}(\mathbb{R}^{d})}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-p)}{2p}}\|\mathcal{F}[f]\|_{H^{\gamma_{1},\gamma_{2}}_{p,in}(\mathbb{R}^{d})}.$
Here $p^{\prime}$ denotes the Hölder conjugate of $p$.
Proof.
First, we prove (i). Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$. By Hölder’s inequality,
$\left|\left\langle\mathcal{F}[f],\varphi\right\rangle\right|$ $=\left|\left\langle(I-\Delta)^{\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f\right),(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\mathcal{F}^{-1}[\varphi]\right)\right\rangle\right|$
$=\left|\int_{\mathbb{R}^{d}}(I-\Delta)^{\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f\right)(x)(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\mathcal{F}^{-1}[\varphi]\right)(x)\mathrm{d}x\right|$
$\leq\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,in}(\mathbb{R}^{d})}\left\|(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\mathcal{F}^{-1}[\varphi]\right)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}.$
Additionally, since $p^{\prime}\in[2,\infty]$, we have
$\left\|(I-\Delta)^{-\gamma_{1}/2}\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}\mathcal{F}^{-1}[\varphi]\right)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d})}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-p)}{2p}}\left\|\left(1+|x|^{2}\right)^{-\gamma_{1}/2}\left(I-\Delta\right)^{-\gamma_{2}/2}\varphi\right\|_{L_{p}(\mathbb{R}^{d},\mathrm{d}x)}.$
Considering $\left(I-\Delta\right)^{\gamma_{2}/2}\left(\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\varphi\right)(x)$ instead of $\varphi(x)$, we have
$\left|\left\langle\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\left(I-\Delta\right)^{\gamma_{2}/2}\mathcal{F}[f],\varphi\right\rangle\right|$ $=\left|\left\langle\mathcal{F}[f],\left(I-\Delta\right)^{\gamma_{2}/2}\left(\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\varphi\right)\right\rangle\right|$
$\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-p)}{2p}}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,in}(\mathbb{R}^{d})}\|\varphi\|_{L_{p}(\mathbb{R}^{d})}.$
Therefore taking the supremum over all $\varphi\in L_{p}(\mathbb{R}^{d})$ so that $\|\varphi\|_{L_{p}(\mathbb{R}^{d})}=1$, we have
$\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{p^{\prime},out}(\mathbb{R}^{d})}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-p)}{2p}}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,in}(\mathbb{R}^{d})}.$
Next, we prove (ii). However, it is an easy consequence of $(i)$ by taking $\mathcal{F}[f]$ instead of $f$ due to Proposition 3.8(i) and the Fourier inversion theorem. ∎
The Fourier and inverse Fourier transforms exhibit particularly favorable behaviors as isometries within the $L_{2}(\mathbb{R}^{d})$-space. This interesting characteristic can also be extended to our weighted Bessel potential spaces. More specifically, when the exponent of integrability $p$ equals 2, the Fourier transform $\mathcal{F}$ acts as an isometry preserving the norms between $H_{2,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$ and $H_{2,out}^{\gamma_{2},\gamma_{1}}(\mathbb{R}^{d})$. To begin, we act the Fourier transform on elements in $H_{2,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$.
Proposition 3.10.
Let $\gamma_{1},\gamma_{2}\in\mathbb{R}$. Then
$f\in H_{2,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{ifand only if}~{}\mathcal{F}[f]\in H_{2,out}^{\gamma_{2},\gamma_{1}}(\mathbb{R}^{d}).$
Additionally,
$\left\|f\right\|_{H_{2,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}=\left\|\mathcal{F}[f]\right\|_{H_{2,out}^{\gamma_{2},\gamma_{1}}(\mathbb{R}^{d})}.$
Proof.
In Proposition 3.9(i), it is already shown that
$\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}\leq\|f\|_{H^{\gamma_{1},\gamma_{2}}_{2,in}(\mathbb{R}^{d})}\quad\forall f\in H^{\gamma_{1},\gamma_{2}}_{2,in}(\mathbb{R}^{d}).$
Thus it suffices to show that
$\|f\|_{H^{\gamma_{1},\gamma_{2}}_{2,in}(\mathbb{R}^{d})}\leq\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}$
for all $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ such that $\mathcal{F}[f]\in H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})$. Let $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ so that $\mathcal{F}[f]\in H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})$. Then
$\left|\left\langle f,\varphi\right\rangle\right|$ $=\left|\left\langle\left(\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}(I-\Delta)^{\gamma_{2}/2}\mathcal{F}[f]\right),\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{1}/2}(I-\Delta)^{-\gamma_{2}/2}\mathcal{F}[\varphi]\right)\right\rangle\right|$
$=\left|\int_{\mathbb{R}^{d}}\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}(I-\Delta)^{\gamma_{2}/2}\mathcal{F}[f](x)\left(1+|\cdot|^{2}\right)^{-\gamma_{1}/2}(I-\Delta)^{-\gamma_{2}/2}\mathcal{F}[\varphi](x)\mathrm{d}x\right|$
$\leq\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}\left\|\left(\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}[\varphi]\right)\right\|_{L_{2}(\mathbb{R}^{d})}$
$=\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}\left\|(I-\Delta)^{-\gamma_{2}/2}\left(1+|\cdot|^{2}\right)^{-\gamma_{1}/2}\varphi\right\|_{L_{2}(\mathbb{R}^{d})},$
where Plancherel’s theorem is used in the last equality. Thus considering $\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}\varphi$ instead of $\varphi$, we have
$\left|\left\langle(I-\Delta)^{\gamma_{1}/2}\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}f,\varphi\right\rangle\right|=\left|\left\langle f,\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}\varphi\right\rangle\right|\leq\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}\left\|\varphi\right\|_{L_{2}(\mathbb{R}^{d})},$
which implies
$\|f\|_{H^{\gamma_{1},\gamma_{2}}_{2,in}(\mathbb{R}^{d})}\leq\|\mathcal{F}[f]\|_{H^{\gamma_{2},\gamma_{1}}_{2,out}(\mathbb{R}^{d})}.$
The proposition is proved. ∎
Likewise, in the case where the integrability exponent $p$ is equal to 2, the Fourier transform $\mathcal{F}$ operates from an outer weighted space to an inner weighted space as an isometry.
Proposition 3.11.
Let $\gamma_{1},\gamma_{2}\in\mathbb{R}$. Then
$f\in H_{2,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{if and only if}~{}\mathcal{F}[f]\in H_{2,in}^{\gamma_{2},\gamma_{1}}(\mathbb{R}^{d}).$
Additionally,
$\left\|f\right\|_{H_{2,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}=\left\|\mathcal{F}[f]\right\|_{H_{2,in}^{\gamma_{2},\gamma_{1}}(\mathbb{R}^{d})}.$
Proof.
It is an easy consequence of Proposition 3.10 by taking $\mathcal{F}[f]$ instead of $f$ with the help of Proposition 3.8 and the Fourier inversion theorem. ∎
Remarkable connections endure between the inner and outer spaces through the Fourier transform even if the integrability exponent $p$ surpasses 2.
Proposition 3.12.
Let $p\in(2,\infty]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, and $\delta\in\left(\frac{d(p-2)}{2(p-1)},\infty\right)$.
  1. (i)
    If $f\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, then
    $\mathcal{F}[f]\in H_{2,in}^{\gamma_{2}-\delta,\gamma_{1}}(\mathbb{R}^{d}).$
    Additionally,
    $\|\mathcal{F}[f]\|_{H_{2,in}^{\gamma_{2}-\delta,\gamma_{1}}(\mathbb{R}^{d})}\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\|f\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
  2. (ii)
    If $\mathcal{F}[f]\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, then
    $f\in H_{2,in}^{\gamma_{2}-\delta,\gamma_{1}}(\mathbb{R}^{d}).$
    Additionally,
    $\|f\|_{H_{2,in}^{\gamma_{2}-\delta,\gamma_{1}}(\mathbb{R}^{d})}\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\|\mathcal{F}[f]\|_{H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}.$
Proof.
It is sufficient to prove (i) since (ii) easily comes from (i) due to Proposition 3.8 and the Fourier inversion theorem. Let $f\in H^{\gamma_{1},\gamma_{2}}_{p,out}(\mathbb{R}^{d})$. Then by Hölder’s inequality,
$\left|\langle\mathcal{F}[f],\varphi\rangle\right|$ $=\left|\left\langle f,\mathcal{F}^{-1}[\varphi]\right\rangle\right|$
$=\left|\left\langle\left(1+|\cdot|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f,\left(1+|\cdot|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi]\right\rangle\right|$
$=\left|\int_{\mathbb{R}^{d}}\left(1+|x|^{2}\right)^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}f(x)\left(1+|x|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi](x)\mathrm{d}x\right|$
$\leq\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,out}(\mathbb{R}^{d})}\left\|\left(1+|x|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi](x)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d},\mathrm{d}x)}.$
Recalling $p^{\prime}\in[1,2)$, we apply Hölder’s inequality again and obtain
$\left\|\left(1+|x|^{2}\right)^{-\gamma_{2}/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi](x)\right\|_{L_{p^{\prime}}(\mathbb{R}^{d},\mathrm{d}x)}$
$\leq\|(1+|\cdot|^{2})^{-\delta/2}\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\left\|(1+|\cdot|^{2})^{(\delta-\gamma_{2})/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi]\right\|_{L_{2}(\mathbb{R}^{d})},$
where $2/0:=\infty$. Since $\delta>\frac{d(p-2)}{2(p-1)}$, we have
$\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}<\infty.$
Thus applying the Plancherel theorem, we have
$\left|\left\langle\mathcal{F}[f],\varphi\right\rangle\right|\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,out}(\mathbb{R}^{d})}\left\|\mathcal{F}\left[(1+|\cdot|^{2})^{(\delta-\gamma_{2})/2}(I-\Delta)^{-\gamma_{1}/2}\mathcal{F}^{-1}[\varphi]\right]\right\|_{L_{2}(\mathbb{R}^{d})}.$
Finally, considering
$\mathcal{F}\left[(I-\Delta)^{\gamma_{1}/2}\left(1+|\cdot|^{2}\right)^{-(\delta-\gamma_{2})/2}\mathcal{F}^{-1}[\varphi]\right]$
$=\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\left(I-\Delta\right)^{-(\delta-\gamma_{2})/2}\varphi$
instead of $\varphi$, we have
$\left|\left\langle\left(I-\Delta\right)^{(\gamma_{2}-\delta)/2}\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\mathcal{F}[f],\varphi\right\rangle\right|$
$=\left|\left\langle\mathcal{F}[f],\left(1+|\cdot|^{2}\right)^{\gamma_{1}/2}\left(I-\Delta\right)^{-(\delta-\gamma_{2})/2}\varphi\right\rangle\right|$
$=\left|\left\langle\mathcal{F}[f],\mathcal{F}\left[(I-\Delta)^{\gamma_{1}/2}\left(1+|\cdot|^{2}\right)^{-(\delta-\gamma_{2})/2}\mathcal{F}^{-1}[\varphi]\right]\right\rangle\right|$
$\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,out}(\mathbb{R}^{d})}\left\|\varphi\right\|_{L_{2}(\mathbb{R}^{d})}.$
Finally, taking the supremum with respect to $\varphi\in\mathcal{S}(\mathbb{R}^{d})$ so that $\|\varphi\|_{L_{2}(\mathbb{R}^{d})}$, we obtain
$\|\mathcal{F}[f]\|_{H_{2,in}^{\gamma_{2}-\delta,\gamma_{1}}}\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta/2}\right\|_{L_{2/(2-p^{\prime})}(\mathbb{R}^{d})}\|f\|_{H^{\gamma_{1},\gamma_{2}}_{p,out}(\mathbb{R}^{d})}.$
The proposition is proved. ∎
We are now ready to establish spaces for handling data within our evolutionary framework, utilizing our newly introduced weighted Bessel potential spaces as a foundation.
We slightly abuse the notation in the following definition as follows : for any $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable function $u$ on $(0,T)$, we put $u(t,\cdot)=u(t)$, i.e.
$\langle u(t,\cdot),\varphi\rangle:=\langle u(t),\varphi\rangle\quad\forall\varphi\in\mathcal{S}(\mathbb{R}^{d}).$
This notation is useful especially when $u(t,\cdot)$ has a realization for each $t\in(0,T)$ by putting $u(t,\xi)=u(t)(\xi)$.
Definition 3.13 (Weighted Bessel potential-valued spaces).
  1. Let $p,q\in[1,\infty]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, and $w(t)$ be a non-negative measurable function on $(0,T)$.
  2. (i)
    We define $\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)$ as the class of all $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable functions $u$ on $(0,T)$ such that
    $\|u\|_{\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)}$ $:=\|u\|_{L_{p,q}\left((0,T),w^{p}(t)\mathrm{d}t;H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})\right)}$
    $:=\int_{0}^{T}\|u(t,\cdot)\|^{p}_{H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})}w^{p}(t)\mathrm{d}t$
    $=\int_{0}^{T}\left\|(I-\Delta)^{\gamma_{1}/2}\left((1+|\cdot|^{2})^{\gamma_{2}/2}u(t,\cdot)\right)\right\|^{p}_{L_{q}}w^{p}(t)\mathrm{d}t<\infty.$
  3. (ii)
    We define $\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)$ as the class of all $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable functions $u$ on $(0,T)$ such that
    $\|u\|_{\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)}$ $:=\|u\|_{L_{p,q}\left((0,T),w^{p}(t)\mathrm{d}t;H^{\gamma_{1},\gamma_{2}}_{q,out}(\mathbb{R}^{d})\right)}$
    $:=\int_{0}^{T}\|u(t,\cdot)\|^{p}_{H^{\gamma_{1},\gamma_{2}}_{q,out}(\mathbb{R}^{d})}w^{p}(t)\mathrm{d}t$
    $=\int_{0}^{T}\left\|(1+|\cdot|^{2})^{\gamma_{2}/2}(I-\Delta)^{\gamma_{1}/2}\left(u(t,\cdot)\right)\right\|^{p}_{L_{q}}w^{p}(t)\mathrm{d}t<\infty.$
  4. (iii)
    We define $\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)$ as the class of all $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable functions $u$ on $(0,T)$ such that
    $\|u\|_{\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in}\left((0,t_{1})\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)}<\infty$
    for all $0<t_{1}<T$.
  5. (iv)
    We define $\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)$ as the class of all $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable functions $u$ on $(0,T)$ such that
    $\|u\|_{\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out}\left((0,t_{1})\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right)}<\infty.$
    for all $0<t_{1}<T$.
  6. (v)
    We define the classes of $\mathcal{S}^{\prime}(\mathbb{R}^{d})$-valued (Borel) measurable functions $u$ on $(0,T)$ whose Fourier transform with respect to the space variable is in weighted Bessel potential spaces. We write
    $u\in\mathcal{F}^{-1}\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    $u\in\mathcal{F}^{-1}\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    $u\in\mathcal{F}^{-1}\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    and
    $u\in\mathcal{F}^{-1}\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    if
    $\mathcal{F}[u(t,\cdot)]\in\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    $\mathcal{F}[u(t,\cdot)]\in\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    $\mathcal{F}[u(t,\cdot)]\in\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    and
    $\mathcal{F}[u(t,\cdot)]\in\mathbf{H}^{\gamma_{1},\gamma_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},w^{p}(t)\mathrm{d}t\right),$
    respectively.
Here we consider the corresponding essential supremum instead of the integration if $p=\infty$. For instance,
$\|u\|_{L_{\infty,q}\left((0,T),w^{p}(t)\mathrm{d}t;H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})\right)}$ $:=\inf\left\{M\in\mathbb{R}:\left|\left\{t\in(0,T):w(t)\|u(t,\cdot)\|^{p}_{H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})}>M\right\}\right|=0\right\}.$
Moreover, we omit $in$ and $out$ if $\gamma_{2}=0$ for simpler notation and identification with the unweighted Bessel potential spaces. $w^{p}(t)\mathrm{d}t$ is also skipped if $w(t)\equiv 1$.
Particularly, inhomogeneous data represented by $f$ to (1.5) could belong to a weighted Bessel potential-valued space, as will be demonstrated in numerous corollaries of Theorem 2.21 in the subsequent sections.
4. Uniqueness of a weak solution with general data
In this section, we establish the uniqueness of a Fourier-space weak solution to (1.5) by introducing a representation of a solution obtained through the use of Sobolev’s mollifier (approximations to the identity). It seems that the space $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ is too expansive to claim the uniqueness of a weak solution. Therefore, our initial step involves constraining the space $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ to a more manageable one where Sobolev’s mollifiers can be effectively applied. Since distributions on $\mathbb{R}^{d}$ include all locally integrable functions on $\mathbb{R}^{d}$, it is obvious that
$\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})\subset\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$
and
$\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\subset L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right).$
Our proof of uniqueness begins by demonstrating the following representation.
Lemma 4.1 (A representation of a solution).
Let $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$, $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$, and $u$ be a Fourier-space weak solution to (1.5). Assume that
$u\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Then
$\mathcal{F}[u(t,\cdot)](\xi)=\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\mathrm{d}s+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s\quad(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d}.$ (4.1)
Proof.
Let $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$. Since $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$, $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$, and $u$ is a Fourier-space weak solution to (1.5) in the space $\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$, we have
$\left\langle\mathcal{F}[u(t,\cdot)],\mathcal{F}[\varphi](\cdot)\right\rangle$ $=\left(\mathcal{F}[u_{0}],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}+\int_{0}^{t}\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s$
$\quad+\int_{0}^{t}\left(\mathcal{F}[f(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$
Here we used realizations mentioned in Remark 2.12. The above equality implies that for almost every $t\in(0,T)$, the distribution $\mathcal{F}[u(t,\cdot)]$ has a realization on $\mathbb{R}^{d}$. More precisely, by using the separability of $\mathcal{D}(\mathbb{R}^{d})$ and Fubini’s theorem, we have
$\left(\mathcal{F}[u(t,\cdot)],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}$ $=\left(\mathcal{F}[u_{0}],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}+\int_{0}^{t}\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s$
$\quad+\int_{0}^{t}\left(\mathcal{F}[f(s,\cdot)],\mathcal{F}[\varphi]\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$ (4.2)
We use Sobolev mollifiers with additional specific properties. Let $\chi$ be a function in $\mathcal{S}(\mathbb{R}^{d})$ so that $\mathcal{F}[\chi]$ is non-negative and symmetric, i.e. $\mathcal{F}[\chi](\xi)=\mathcal{F}[\chi](-\xi)\geq 0$ for all $\xi\in\mathbb{R}^{d}$. Additionally, assume that $\mathcal{F}[\chi]$ has a compact support and
$\int_{\mathbb{R}^{d}}\chi(x)\mathrm{d}x=(2\pi)^{d/2}.$
For $\varepsilon\in(0,1)$, denote
$\chi^{\varepsilon}(x):=\frac{1}{\varepsilon^{d}}\chi\left(\frac{x}{\varepsilon}\right).$
Fix $x\in\mathbb{R}^{d}$ and put $\chi^{\varepsilon}(x-\cdot)$ in (4.2) instead of $\varphi$. Then
$\int_{\mathbb{R}^{d}}\mathcal{F}[u(t,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\mathcal{F}[u_{0}](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi+\int_{0}^{t}\int_{\mathbb{R}^{d}}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi\mathrm{d}s$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathcal{F}[f(s,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$
Thus recalling properties of $\mathcal{F}[\chi]$ and the Fourier transform, we have
$\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}\mathcal{F}[u(t,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\mathcal{F}[u(t,\cdot)](\xi)\overline{\mathrm{e}^{-ix\cdot\xi}\mathcal{F}[\chi^{\varepsilon}](-\xi)}\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\mathcal{F}[u(t,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\mathcal{F}[u_{0}](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi+\int_{0}^{t}\int_{\mathbb{R}^{d}}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi\mathrm{d}s$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}}\mathcal{F}[f(s,\cdot)](\xi)\overline{\mathcal{F}[\chi^{\varepsilon}(x-\cdot)](\xi)}\mathrm{d}\xi\mathrm{d}s$
$=\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}\mathcal{F}[u_{0}](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}\xi+\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}\left(\int_{0}^{t}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}s\right)\mathrm{d}\xi$
$\quad+\int_{\mathbb{R}^{d}}\mathrm{e}^{ix\cdot\xi}\left(\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}s\right)\mathrm{d}\xi\quad(a.e.)~{}t\in(0,T).$
Here Fubini’s theorem could be applicable to all terms above since $\mathcal{F}[\chi^{\varepsilon}]\in\mathcal{D}(\mathbb{R}^{d})$, $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$, $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$, and
$u\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Due to the Fourier inversion theorem,
$\mathcal{F}[u(t,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)=\mathcal{F}[u_{0}](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)+\int_{0}^{t}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}s+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathcal{F}[\chi^{\varepsilon}](\xi)\mathrm{d}s$ (4.3)
for almost every $(t,\xi)\in(0,T)\times\mathbb{R}^{d}$. Observe that
$\mathcal{F}[\chi^{\varepsilon}](\xi)=\mathcal{F}[\chi](\varepsilon\xi)$
and
$\mathcal{F}[\chi](0)=(2\pi)^{-d/2}\int_{\mathbb{R}^{d}}\chi(y)\mathrm{d}y=1.$
Finally, taking $\varepsilon\downarrow 0$ in (4.3), we have (4.1). ∎
Remark 4.2.
It is easy to show that (4.1) with the Fubini theorem implies
$u(t,x)=u_{0}(x)+\int_{0}^{t}\psi(s,-i\nabla)u(s,\cdot)(x)\mathrm{d}s+\int_{0}^{t}f(s,x)\mathrm{d}s\quad(a.e.)~{}t\in(0,T).$
This equation should be interpreted as elements within $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, given that the term $\psi(s,-i\nabla)u(s,\cdot)(x)$ is understood as the inverse Fourier transform of a locally integrable function $\psi(s,\xi)\mathcal{F}u(s,\cdot)$ with respect to $\xi$, as discussed in Remark 2.12.
We can now demonstrate the uniqueness of the solution using the provided representation and taking advantage of the linearity of (1.5). It is obvious that (4.1) implies
$u\in\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right).$ (4.4)
This property heavily depends on conditions of data $u_{0}$ and $f$ since they are important used in the representation formula. However, these restrictions on the initial data $u_{0}$ and the inhomogeneous data $f$ can be eliminated due to the linearity if we merely have interest in the uniqueness. In particular, we can claim the uniqueness of a solution even for $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$-valued data.
Theorem 4.3 (Uniquness of a weak solution).
Let $u_{0}\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ and $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$. Then a Fourier-space weak solution to (1.5) is unique in
$\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Proof.
Let $u_{1}$ and $u_{2}$ be Fourier-space weak solutions to (1.5) so that
$u_{1},u_{2}\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Put
$u=u_{1}-u_{2}.$
Then by Lemma 4.1, we have
$\mathcal{F}[u(t,\cdot)](\xi)=\int_{0}^{t}\psi(s,\xi)\mathcal{F}[u(s,\cdot)](\xi)\mathrm{d}s\quad(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d}.$
Due to Grönwall’s inequality, we show that
$\mathcal{F}[u(t,\cdot)](\xi)=0\quad(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d}.$
Thus
$\mathcal{F}[u_{1}(t,\cdot)](\xi)=\mathcal{F}[u_{2}(t,\cdot)](\xi)\quad(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d},$ (4.5)
which implies that $u_{1}=u_{2}$ as an element in
$\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Remark 4.4.
The uniqueness stated in Theorem 4.3 may appear uncertain when considering a solution $u$ for (1.5) as a function valued in $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, defined over the interval $(0,T)$ if $\psi(t,\xi)$ vanishes at certain points as explained in Remark 2.13. Additionally, we do not know if $u_{1}$ and $u_{2}$ satisfy (4.4) in the proof of Theorem 4.3 since the conditions on $u_{0}$ and $f$ are weakened. Nevertheless, we have shown that the realizations $\mathcal{F}[u_{1}(t,\cdot)](\xi)$ and $\mathcal{F}[u_{2}(t,\cdot)](\xi)$ coincide for almost every $(t,\xi)\in(0,T)\times\mathbb{R}^{d}$, considering all solutions $u_{1}$ and $u_{2}$ to the equation (1.5) in
$\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right),$
[…]
Moreover, this issue does not cause any problem in our main theorem, Theorem 2.21, because we establish both the existence and uniqueness of a solution $u$ within the intersection of two function spaces $\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ and $\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$. More precisely,
$u\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$
implies
$u\in\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$
since $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ as mentioned (4.4). In other words, the ambiguity of the uniqueness of $u$ as a $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$-valued function on $(0,T)$ could be resolved completely since $u$ is also an element of $\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ in Theorem 2.21.
Next, we characterize conditions on symbols $\psi(t,\xi)$ to preserve the uniqueness of a solution whose Fourier transform is in a weighted $L_{p,q}$-space. This is readily accomplished due to Hölder’s inequality.
Corollary 4.5.
Let $p,q\in[1,\infty]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, $u_{0}\in\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$, $f\in L_{1,t\text{-}loc}\left((0,T);\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})\right)$, and $W(t,x)$ be a positive measurable function on $(0,T)\times\mathbb{R}^{d}$. Assume that
$\psi\in L_{p^{\prime},q^{\prime},t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1/W(t,\xi))^{q^{\prime}}\mathrm{d}\xi\right),$
where $p^{\prime}$ and $q^{\prime}$ are Hölder conjugates of $p$ and $q$, respectively. Then a Fourier-space weak solution to (1.5) is unique in
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(W(t,\xi))^{q}\mathrm{d}\xi\right).$
Proof.
Let
$u\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(W(t,\xi))^{q}\mathrm{d}\xi\right)$
be a Fourier-space weak solution to (1.5). It is sufficient to show that
$u\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$
due to Theorem 4.3. However, it is an easy consequence of Hölder’s inequality. Indeed, for any $t\in(0,T)$ and $R\in(0,\infty)$, we have
$\int_{0}^{t}\int_{B_{R}}|\psi(s,\xi)||\mathcal{F}[u](\xi)|\mathrm{d}\xi\mathrm{d}s$ $=\int_{0}^{t}\int_{B_{R}}|\psi(s,\xi)|\frac{1}{W(t,\xi)}|\mathcal{F}[u](\xi)|W(t,\xi)\mathrm{d}\xi\mathrm{d}s$
$\leq\|\psi\|_{L_{p^{\prime},q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}t(1/W(t,\xi))^{q^{\prime}}\mathrm{d}\xi\right)}\|\mathcal{F}[u]\|_{L_{p,q}\left((0,t)\times B_{R},\mathrm{d}t(W(t,\xi))^{q}\mathrm{d}\xi\right)}<\infty.$
Take note that in the aforementioned corollary, the weight denoted as $W$ is assumed to be positive. This requirement is maintained in order to ensure the existence of a solution in our theory, as seen in Theorem 2.34. However, there is a possibility to relax this positive constraint on weights by considering a more intricate limiting process, which seems to be an interesting topic for further investigations (cf. [19, 21]).
5. Proof and Corollaries of Theorem 2.21
We are now ready to initiate the proof of our main theorem. It is important to recall that the uniqueness has already been established due to Theorem 4.3. Thus our current goal becomes to show the existence of a Fourier-space weak solution $u$ to (1.5). A credible candidate for this solution $u$ is presented in (2.12). Therefore, our task is to verify that the function $u$ as defined in (2.12) indeed satisfies the criteria for being our weak solution in the forthcoming proof.
Proof of Theorem 2.21.
First, observe that the mapping
$(t,\xi)\mapsto\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
is in $L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$ due to (2.9). Here we used the continuity of the functions
$t\mapsto\int_{B_{R}}\exp\left(\int_{0}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi$
and
$t\mapsto\int_{B_{R}}\int_{0}^{t}\exp\left(\int_{s}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi,$
i.e. for any $t\in(0,T)$, there exist $t_{0},t_{1}\in[0,t]$ so that
$\int_{B_{R}}\exp\left(\int_{0}^{t_{0}}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi=\sup_{\rho\in[0,t]}\int_{B_{R}}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi$
and
$\int_{B_{R}}\int_{0}^{t_{1}}\exp\left(\int_{s}^{t_{1}}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$
$=\sup_{\rho\in[0,t]}\int_{B_{R}}\int_{0}^{\rho}\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi.$
Similarly, the mapping
$(t,\xi)\mapsto\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\xi)]\mathrm{d}s$
is in
$L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right)$
due to (2.10). Thus recalling that
$u(t,x):=\mathcal{F}^{-1}\left[\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]\right](x)+\mathcal{F}^{-1}\left[\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s\right](x),$
we have
$u\in\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
[…]
$\mathcal{F}[u(t,\cdot)](\xi)$
$=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\int_{s}^{t}\frac{\mathrm{d}}{\mathrm{d}\rho}\left[\exp\left(\int_{s}^{\rho}\psi(r,\xi)\mathrm{d}r\right)\right]\mathrm{d}\rho\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\int_{s}^{t}\left[\psi(\rho,\xi)\exp\left(\int_{s}^{\rho}\psi(r,\xi)\mathrm{d}r\right)\right]\mathrm{d}\rho\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$\quad+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\psi(\rho,\xi)\int_{0}^{\rho}\exp\left(\int_{s}^{\rho}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s\mathrm{d}\rho+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)-\int_{0}^{t}\psi(\rho,\xi)\exp\left(\int_{0}^{\rho}\psi(r,\xi)\mathrm{d}r\right)\mathrm{d}\rho\mathcal{F}[u_{0}](\xi)$
$\quad+\int_{0}^{t}\psi(\rho,\xi)\mathcal{F}[u(\rho,\cdot)](\xi)\mathrm{d}\rho+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$
$=\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\psi(\rho,\xi)\mathcal{F}[u(\rho,\cdot)](\xi)\mathrm{d}\rho+\int_{0}^{t}\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$ (5.1)
for almost every $(t,\xi)\in(0,T)\times\mathbb{R}^{d}$. The condition (2.10) is used above to apply the Fubini theorem. Finally, for any $\varphi\in\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d})$, taking the integration with $\overline{\mathcal{F}[\varphi]}$ in (5.1) and applying the definitions of actions on the elements of the class $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$ with the Fubini theorem, we have
$\left\langle u(t,\cdot),\varphi\right\rangle=\langle u_{0},\varphi\rangle+\int_{0}^{t}\left(\psi(s,\cdot)\mathcal{F}[u(s,\cdot)],\mathcal{F}[\varphi](\cdot)\right)_{L_{2}(\mathbb{R}^{d})}\mathrm{d}s+\int_{0}^{t}\left\langle f(s,\cdot),\varphi\right\rangle\mathrm{d}s\quad(a.e.)~{}t\in(0,T),$
which implies that $u$ is a Fourier-space weak solution (1.5) as discussed in Remark 2.12. ∎
Now, let’s examine data associated with general weights. We present a series of corollaries to demonstrate that our mathematical requirements for $\psi$ are generous enough to accommodate a wide range of weighted data. It is important to note that we are exclusively addressing positive weights.
Corollary 5.1.
Let $q\in[1,\infty]$, $W_{0}(\xi)$ and $W_{1}(\xi)$ be positive measurable functions on $\mathbb{R}^{d}$, and $W_{2}(t,\xi)$ be a positive measurable function on $(0,T)\times\mathbb{R}^{d}$ so that both $W_{0}(\xi)$ and $W_{1}(\xi)$ are locally bounded below and $W_{1}(\xi)$ is a lower bound of $W_{2}(t,\xi)$, i.e. for each $t\in(0,T)$ and $R\in(0,\infty)$, there exist positive constants $\kappa_{0}(R)$ and $\kappa_{1}(R)$ so that
$W_{0}(\xi)\geq\kappa_{0}(R)\quad\forall\xi\in B_{R}.$ (5.2)
and
$W_{2}(s,\xi)\geq W_{1}(\xi)\geq\kappa_{1}(R)\quad\forall(s,\xi)\in(0,t)\times B_{R}.$ (5.3)
Assume that $u_{0}\in\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d},W_{0}^{q}(\xi)\mathrm{d}\xi)$ and $f\in\mathcal{F}^{-1}L_{1,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tW_{2}^{q}(t,\xi)\mathrm{d}\xi\right)$. Additionally, suppose that the symbol $\psi(t,\xi)$ satisfies the following integrability conditions
$\int_{0}^{t}\left\|\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{q^{\prime}}\left(B_{R},\mathrm{d}\xi\right)}\mathrm{d}\rho$
$+\int_{0}^{t}\left\|\frac{1+|\psi(\rho,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)|]\right)\mathrm{d}r\right\|_{L_{q^{\prime}}\left(B_{R},\mathrm{d}\xi\right)}\mathrm{d}\rho<\infty$ (5.4)
for all $t\in(0,T)$ and $R\in(0,\infty)$. Then there exists a unique Fourier-space weak solution to (1.5) in
$\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Proof.
We use Theorem 2.21. First of all, it is easy to check that $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and
$f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$
due to Hölder’s inequality and the local lower bounds of $W_{0}$ and $W_{2}$ given in (5.2) and (5.3). Next we show that (5.4) implies (2.9) and (2.10). Recall that $u_{0}\in\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d},W_{0}^{q}(\xi)\mathrm{d}\xi)$ and
$f\in L_{1,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tW_{2}^{q}(t,\xi)\mathrm{d}\xi\right).$
Thus by Hölder’s inequality and (5.4),
$\int_{B_{R}}\int_{0}^{t}(1+|\psi(\rho,\xi)|)\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\rho\mathrm{d}\xi$
$=\int_{0}^{t}\int_{B_{R}}\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)W_{0}(\xi)\mathrm{d}\xi\mathrm{d}\rho$
$\leq\int_{0}^{t}\left\|\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\right\|_{L_{q^{\prime}}\left(B_{R},\mathrm{d}\xi\right)}\mathrm{d}\rho\|\mathcal{F}[u_{0}]\|_{L_{q}\left(B_{R},\mathrm{d}sW_{0}^{q}(\xi)\mathrm{d}\xi\right)}<\infty$
for all $t\in(0,T)$ and $R\in(0,\infty)$. Similarly, by the Fubini theorem, the Hölder inequality, (5.3), and (5.4),
$\int_{B_{R}}\int_{0}^{t}(1+|\psi(\rho,\xi)|)\int_{0}^{\rho}\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\rho\mathrm{d}\xi$
$=\int_{0}^{t}\int_{0}^{\rho}\int_{B_{R}}\frac{1+|\psi(\rho,\xi)|}{W_{2}(s,\xi)}\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|W_{2}(s,\xi)\mathrm{d}\xi\mathrm{d}s\mathrm{d}\rho$
$\leq\int_{0}^{t}\int_{0}^{t}\int_{B_{R}}\frac{1+|\psi(\rho,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)W_{2}(s,\xi)\right|\mathrm{d}\xi\mathrm{d}s\mathrm{d}\rho$
$\leq\left\|(1+|\psi(\rho,\xi)|)\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{1,q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}\rho(1/W_{1}(\xi))^{q^{\prime}}\mathrm{d}\xi\right)}\|\mathcal{F}[f(s,\cdot)(\xi)\|_{L_{1,q}\left((0,t)\times B_{R},\mathrm{d}sW_{2}^{q}(s,\xi)\mathrm{d}\xi\right)}$
<$\infty$< ∞
for all $t\in(0,T)$ and $R\in(0,\infty)$. Therefor by Theorem 2.21, there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right).$
Remark 5.2.
We used a very rough estimate to control the term
$\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\leq\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)$
in the proof of Corollary 5.1 since there is a possibility that the real part of $\psi(r,\xi)$ is non-negative for all $r$ and $\xi$. However, this rough estimate makes us lose all decay of the exponential term derived from negative values of the real part of the symbol $\psi(s,\xi)$ in general. On the other hand, recall that the conditions (2.9) and (2.10) in Theorem 2.21 do not require controls of the term $\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)$. Especially, Theorem 2.21 works for symbols that are not locally bounded by confirming conditions (2.9) and (2.10). In other words, our main theorem could apply to more general sign changing symbols by estimating the term $\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)$ for all $s<\rho$ instead of using the rough estimate above. This kind of example can be found in the proof of Theorem 2.27 (see Section 6 below).
If our input data $f$ exhibits better integrability concerning the time variable, it may lead to the solution $u$ being confined to a more restricted space, thereby demonstrating enhanced integrability in relation to the time variable. This can be accomplished through local estimates that are readily derived thanks to Hölder’s inequality.
Corollary 5.3.
Let $p,q\in[1,\infty]$, $W_{0}(\xi)$ and $W_{1}(\xi)$ be positive measurable functions defined on $\mathbb{R}^{d}$, and $W_{2}(t,\xi)$ be a positive measurable function on $(0,T)\times\mathbb{R}^{d}$ so that both $W_{0}(\xi)$ and $W_{1}(\xi)$ are locally bounded below and $W_{1}(\xi)$ is a lower bound of $W_{2}(t,\xi)$, i.e. for each $t\in(0,T)$ and $R\in(0,\infty)$, there exist positive constants $\kappa_{0}(R)$ and $\kappa_{1}(R)$ so that
$W_{0}(\xi)\geq\kappa_{0}(R)\quad\forall\xi\in B_{R}$
and
$W_{2}(s,\xi)\geq W_{1}(\xi)\geq\kappa_{1}(R)\quad\forall(s,\xi)\in(0,t)\times B_{R}.$ (5.5)
Assume that $u_{0}\in\mathcal{F}^{-1}L_{q,\ell oc}(\mathbb{R}^{d},W_{0}^{q}(\xi)\mathrm{d}\xi)$ and $f\in\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tW_{2}^{q}(t,\xi)\mathrm{d}\xi\right)$. Additionally, suppose that the symbol $\psi(t,\xi)$ satisfies the following condition:
$\left\|\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
$+\left\|\frac{1+|\psi(\rho,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[\psi(r,\xi)]\right|\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$ (5.6)
for all $t\in(0,T)$ and $R\in(0,\infty)$. Then there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{p,q,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1+|\psi(t,\xi)|)\mathrm{d}\xi\right).$
Moreover, for each $t\in(0,T)$ and $R\in(0,\infty)$, the solution $u$ satisfies
$\left\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\right\|_{L_{p,q}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}$
$\leq\left\|\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|\mathcal{F}[u_{0}]\|_{L_{q}\left(B_{R},\mathrm{d}sW_{0}^{q}(\xi)\mathrm{d}\xi\right)}$
$\quad+t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|\mathcal{F}[f(s,\cdot)(\xi)\|_{L_{p,q}\left((0,t)\times B_{R},\mathrm{d}sW_{2}^{q}(\xi)\mathrm{d}\xi\right)},$ (5.7)
where $p^{\prime}$ is the Hölder conjugate of $p$ and $1/\infty:=0$.
Proof.
Observe that (5.6) implies (5.4) due to Hölder’s inequality. Thus by Corollary 5.1, there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t|\psi(t,\xi)|\mathrm{d}\xi\right),$
which is given by
$u(t,x)=\mathcal{F}^{-1}\left[\exp\left(\int_{0}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[u_{0}]\right](x)+\mathcal{F}^{-1}\left[\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\cdot)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)]\mathrm{d}s\right](x)$
in the sense of (2.11). Thus taking the Fourier transforms above, we have
$\mathcal{F}[u(t,\cdot)](\xi)=\exp\left(\int_{0}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)+\int_{0}^{t}\exp\left(\int_{s}^{t}\psi(r,\xi)\mathrm{d}r\right)\mathcal{F}[f(s,\cdot)](\xi)\mathrm{d}s$ (5.8)
$(a.e.)~{}(t,\xi)\in(0,T)\times\mathbb{R}^{d}$. Note that the last two terms in (5.7) are finite due to assumptions on weights and the conditions on $u_{0}$ and $f$. It suffices to show (5.7). By (5.8), the Hölder inequality, the generalized Minkowski inequality, and (5.5),
$\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\|_{L_{q}(B_{R})}$
$\leq\left\|\frac{1+|\psi(s,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{s}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{\infty}\left(B_{R},\mathrm{d}\xi\right)}\|\mathcal{F}[u_{0}]\|_{L_{q}\left(B_{R},W_{0}^{q}(\xi)\mathrm{d}\xi\right)}$
$\quad+\left\|\frac{1+|\psi(s,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{s}|\Re[\psi(r,\xi)]|\right)\mathrm{d}r\right\|_{L_{\infty}\left(B_{R},\mathrm{d}\xi\right)}\int_{0}^{s}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{q}\left(B_{R},W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho$
$(a.e.)~{}s\in(0,T)$. Therefore taking the $L_{p}$-norms with respect to the variable $s$, we have
$\left\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\right\|_{L_{p,q}\left((0,t)\times B_{R},\mathrm{d}\xi\right)}$
$\leq\left\|\frac{1+|\psi(s,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{s}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}\|\mathcal{F}[u_{0}]\|_{L_{q}\left(B_{R},\mathrm{d}sW_{0}^{q}(\xi)\mathrm{d}\xi\right)}$
$\quad+\left\|\frac{1+|\psi(s,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{s}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}\sup_{s\in(0,t)}\int_{0}^{s}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{q}\left(B_{R},W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho.$ (5.9)
Note that for any $t\in(0,T)$,
$\sup_{s\in(0,t)}\int_{0}^{s}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{q}\left(B_{R},W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho$ $\leq\int_{0}^{t}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{q}\left(B_{R},W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho$
$\leq t^{1/p^{\prime}}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{p,q}\left((0,t)\times B_{R},\mathrm{d}\rho W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho.$ (5.10)
Combining (5.9) and (5.10), we finally have (5.7). The corollary is proved. ∎
At last, we are fully equipped to establish a link between the two weighted Bessel potential spaces introduced in Section 3 and our theories regarding well-posedness.
Recall that the exponents of weighted spaces play crucial roles in applying embedding inequalities. Additionally, the weighted Bessel potential spaces become identical with classical Bessel potential spaces if the weighted exponents are zero. We begin by presenting a corollary to show that data within the classical Bessel potential spaces can be included in our well-posedness theories when the integral exponents with respect to the spatial variable are constrained within the range of $[1,2]$.
Corollary 5.4.
Let $p\in[1,\infty]$, $q\in[1,2]$, $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $u_{0}\in H_{q}^{\gamma_{1}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1}}_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Assume that the symbol $\psi(t,\xi)$ satisfies
$\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$ (5.11)
for all $t\in(0,T)$ and $R\in(0,\infty)$, where $\gamma=\min\{\gamma_{1},\tilde{\gamma}_{1}\}$. Then there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{p,q^{\prime},t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1+|\psi(t,\xi)|)\mathrm{d}\xi\right).$
Moreover, for any $t\in(0,T)$ and $R\in(0,\infty)$, the solution $u$ satisfies
$\left\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\right\|_{L_{p,q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}$
$\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|u_{0}\|_{H^{\gamma_{1}}_{q}(\mathbb{R}^{d})}$
$\quad+\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1}}_{p,q,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)},$
where $p^{\prime}$ is the Hölder conjugate of $p$ and $1/\infty:=0$.
Proof.
Put
$W_{0}(\xi)=W_{1}(\xi)=W_{2}(t,\xi)=\left(1+|\xi|^{2}\right)^{\gamma/2}$
for all $t$ and $\xi$. Due to (5.11), it is obvious that
$\left\|\frac{1+|\psi(\rho,\xi)|}{W_{0}(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
$+\left\|\frac{1+|\psi(\rho,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[\psi(r,\xi)]\right|\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty.$
By proposition 3.9, $u_{0}\in\mathcal{F}^{-1}H_{q^{\prime},out}^{0,\gamma_{1}}(\mathbb{R}^{d})$ and $f\in\mathcal{F}^{-1}\mathbf{H}^{0,\tilde{\gamma}_{1}}_{p,q^{\prime},out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t\mathrm{d}x\right)$. Additionally,
$\|\mathcal{F}[u_{0}]\|_{H^{0,\gamma_{1}}_{q^{\prime},out}(\mathbb{R}^{d})}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}\|u_{0}\|_{H^{\gamma_{1}}_{q}(\mathbb{R}^{d})}$
and
$\|\mathcal{F}[f]\|_{\mathbf{H}^{0,\tilde{\gamma}_{1}}_{p,q^{\prime},out,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1}}_{p,q,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}.$
Thus it is easy to show that for all $t\in(0,T)$ and $R\in(0,\infty)$,
$\|\mathcal{F}[u_{0}]\|_{L_{q^{\prime}}(B_{R},W^{q^{\prime}}(\xi)\mathrm{d}\xi)}\leq\|\mathcal{F}[u_{0}]\|_{L_{q^{\prime}}(\mathbb{R}^{d},W^{q^{\prime}}(\xi)\mathrm{d}\xi)}$ $\leq\|\mathcal{F}[u_{0}]\|_{H^{0,\gamma_{1}}_{q^{\prime},out}(\mathbb{R}^{d})}$
$\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}\|u_{0}\|_{H^{\gamma_{1}}_{q}(\mathbb{R}^{d})}$
and
$\|\mathcal{F}[f]\|_{L_{p,q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}sW^{q^{\prime}}(\xi)\mathrm{d}\xi\right)}\leq\|\mathcal{F}[f]\|_{L_{p,q^{\prime}}\left((0,t)\times\mathbb{R}^{d},\mathrm{d}sW^{q^{\prime}}(\xi)\mathrm{d}\xi\right)}$ $\leq\|\mathcal{F}[f]\|_{\mathbf{H}^{0,\tilde{\gamma}_{1}}_{p,q^{\prime},out,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}$
$\leq\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1}}_{p,q,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}.$
Thus the conclusion directly comes from Corollary 5.3. ∎
Remark 5.5.
The restriction on the integral exponent $q\in[1,2]$ is not crucial in general if the solution $u$ has a nice stability with respect to data $u_{0}$ and $f$ by applying a standard duality argument (cf. [15, 16]).
Next, we examine data within inner weighted spaces featuring non-zero weighted exponents. Unfortunately, this imposes specific constraints on the range of exponents, which arise as a natural consequence of the embedding inequalities.
Corollary 5.6.
Let $p\in[1,\infty]$, $q\in(1,2]$, $\gamma_{1},\tilde{\gamma}_{1}\in\left(-\frac{d(p-1)}{p},\infty\right)$, $\gamma_{2},\tilde{\gamma}_{2}\in[0,\infty)$, $u_{0}\in H_{q,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Assume that the symbol $\psi(t,\xi)$ satisfies
$\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$
for all $t\in(0,T)$ and $R\in(0,\infty)$, where $\gamma=\min\{\gamma_{1},\tilde{\gamma}_{1}\}$. Then there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{p,q^{\prime},t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1+|\psi(t,\xi)|)\mathrm{d}\xi\right).$
Moreover, for any $t\in(0,T)$ and $R\in(0,\infty)$, the solution $u$ satisfies
$\left\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\right\|_{L_{p,q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}$
$\lesssim\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|u_{0}\|_{H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})}$
$\quad+t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)},$
where $p^{\prime}$ is the Hölder conjugate of $p$ and $1/\infty:=0$.
Proof.
The proof of this corollary is almost identical with that of Corollary 5.4. We only highlight the major difference. Due to Proposition 3.4, we may assume that
$\gamma_{1},\tilde{\gamma}_{1}\in(-d(p-1)/p,d/p)$
without loss of generality. Moreover, Proposition 3.7 is used to control $L_{q^{\prime}}$-norms of the Fourier transforms of data in order to consider general regularity exponents of outer spaces instead of only 0. Indeed, for all $t\in(0,T)$ and $R\in(0,\infty)$,
$\|\mathcal{F}[u_{0}]\|_{L_{q^{\prime}}(B_{R},(1+|\xi|^{2})^{(q^{\prime}\gamma)/2}\mathrm{d}\xi)}\leq\|\mathcal{F}[u_{0}]\|_{L_{q^{\prime}}(\mathbb{R}^{d},(1+|\xi|^{2})^{(q^{\prime}\gamma)/2}\mathrm{d}\xi)}\leq\|\mathcal{F}[u_{0}]\|_{H^{0,\gamma_{1}}_{q^{\prime},out}(\mathbb{R}^{d})}$ $\lesssim\|\mathcal{F}[u_{0}]\|_{H^{\gamma_{2},\gamma_{1}}_{q^{\prime},out}(\mathbb{R}^{d})}$
$\lesssim\|u_{0}\|_{H^{\gamma_{1},\gamma_{2}}_{q,in}(\mathbb{R}^{d})}$
and
$\|\mathcal{F}[f]\|_{L_{p,q^{\prime}}\left((0,t)\times B_{R},\mathrm{d}s(1+|\xi|^{2})^{(q^{\prime}\gamma)/2}\mathrm{d}\xi\right)}$ $\leq\|\mathcal{F}[f]\|_{L_{p,q^{\prime}}\left((0,t)\times\mathbb{R}^{d},\mathrm{d}s(1+|\xi|^{2})^{(q^{\prime}\gamma)/2}\mathrm{d}\xi\right)}$
$\leq\|\mathcal{F}[f]\|_{\mathbf{H}^{0,\tilde{\gamma}_{1}}_{p,q^{\prime},out,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}$
$\lesssim\|\mathcal{F}[f]\|_{\mathbf{H}^{\tilde{\gamma}_{2},\tilde{\gamma}_{1}}_{p,q^{\prime},out,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}$
$\lesssim\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)}.$
The corollary is proved. ∎
Moving forward, our focus shifts to the parameter $q$ which exceeds the value of 2. In simpler terms, we are now capable of dealing with the data $u_{0}$ and $f$ in the outer weighted spaces, thanks to the embeddings developed in Section 3, even when $q$ is within the range of $(2,\infty]$.
Corollary 5.7.
Let $p\in[1,\infty]$, $q\in(2,\infty]$, $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $\gamma_{2},\tilde{\gamma}_{2}\in\left(\frac{d(q-2)}{2(q-1)},\infty\right)$, $u_{0}\in H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Assume that the symbol $\psi(t,\xi)$ satisfies
$\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$ (5.12)
for all $t\in(0,T)$ and $R\in(0,\infty)$, where $\gamma=\min\{\gamma_{1},\tilde{\gamma}_{1}\}$. Then there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{p,2,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1+|\psi(t,\xi)|)\mathrm{d}\xi\right).$
Moreover, for any $t\in(0,T)$ and $R\in(0,\infty)$, the solution $u$ satisfies
$\left\|(1+|\psi(s,\xi)|)\mathcal{F}[u(s,\cdot)](\xi)\right\|_{L_{p,2}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}$
$\lesssim_{d,p,q,\gamma_{2},\tilde{\gamma}_{2}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|u_{0}\|_{H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}$
$\quad+t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out}\left((0,t)\times\mathbb{R}^{d}\right)},$ (5.13)
where $p^{\prime}$ is the Hölder conjugate of $p$ and $1/\infty:=0$.
Proof.
It is another easy application of Corollary 5.3. First, we apply an embedding inequality developed in Proposition 3.12. Put
$\delta_{1}=\gamma_{2}~{}\text{and}~{}\delta_{2}=\tilde{\gamma}_{2}.$
Then by Proposition 3.12, we have
$\|\mathcal{F}[u_{0}]\|_{H_{2,in}^{0,\gamma_{1}}(\mathbb{R}^{d})}\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta_{1}/2}\right\|_{L_{2/(2-q^{\prime})}(\mathbb{R}^{d})}\|u_{0}\|_{H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})}$ (5.14)
and
$\|\mathcal{F}[f](t,\cdot)\|_{H_{2,in}^{0,\tilde{\gamma}_{1}}(\mathbb{R}^{d})}\leq\left\|\left(1+|\cdot|^{2}\right)^{-\delta_{2}/2}\right\|_{L_{2/(2-q^{\prime})}(\mathbb{R}^{d})}\|f(t,\cdot)\|_{H_{q,out}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}(\mathbb{R}^{d})}\quad\forall t\in(0,T).$ (5.15)
In particular, $u_{0}\in\mathcal{F}^{-1}L_{2,\ell oc}(\mathbb{R}^{d},W^{2}(\xi)\mathrm{d}\xi)$ and $f\in\mathcal{F}^{-1}L_{p,2,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}tW^{2}(\xi)\mathrm{d}\xi\right)$ with $W(\xi)=(1+|\xi|^{2})^{\gamma/2}$. Additionally, due to (5.12), we have
$\left\|\frac{1+|\psi(\rho,\xi)|}{W(\xi)}\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\right)\mathrm{d}r\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}<\infty$
for all $t\in(0,T)$ and $R\in(0,\infty)$. Thus by applying Corollary 5.3 with
$W(\xi)=W_{0}(\xi)=W_{1}(\xi)=W_{2}(t,\xi),$
there exists a unique Fourier-space weak solution $u$ to (1.5) in
$\mathcal{F}^{-1}L_{p,2,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t(1+|\psi(t,\xi)|)\mathrm{d}\xi\right)$
and this solution $u$ satisfies (5.13) due to (5.7), (5.14), and (5.15). ∎
6. Logarithmic operators with complex-valued coefficients and proof of Theorem 2.27
Demonstrating the well-posedness of evolutionary equations involving the logarithmic Laplacian is an important application of our work. We have already proposed natural extensions of the operator with coefficients in (2.13). We begin this section by providing a proof for Theorem 2.27.
Proof of Theorem 2.27.
Put
$\psi(t,\xi)=\beta(t)\log\left(\psi_{exp}(t,\xi)\right)$
and we show that the symbol $\psi(t,\xi)$ and data $u_{0}$ and $f$ satisfy (2.9) and (2.10). Observe that
$\Re[\psi(t,\xi)]=\Re[\beta(t)]\log\left(|\psi_{exp}(t,\xi)|\right)-\Im[\beta(t)]\arg[\psi_{exp}(t,\xi)]$
and recall that the real-valued exponential function is convex on any open interval of $\mathbb{R}$, where $\arg[\psi_{exp}(t,\xi)]$ denotes the argument of the complex number $\psi_{exp}(t,\xi)$. Thus for all $t\in(0,T)$ and $R\in(0,\infty)$, applying Jensen’s inequality and considering the standard branch cut, we have
$\int_{B_{R}}\exp\left(\int_{0}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi+\int_{B_{R}}\int_{0}^{t}\exp\left(\int_{s}^{t}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$
$\leq\int_{B_{R}}\exp\left(\frac{1}{t}\int_{0}^{t}\log\left(|\psi_{exp}(r,\xi)|^{t\Re[\beta(r)]}\right)\mathrm{d}r+2\pi\int_{0}^{t}|\Im[\beta(r)]|\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi$
$\quad+\int_{B_{R}}\int_{0}^{t}\exp\left(\frac{1}{t-s}\int_{s}^{t}\log\left[\left(|\psi_{exp}(r,\xi)|^{(t-s)\Re[\beta(r)]}\right)\right]\mathrm{d}r+2\pi\int_{0}^{t}|\Im[\beta(r)]|\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi$
$\leq\exp\left(2\pi\int_{0}^{t}|\Im[\beta(r)]|\mathrm{d}r\right)\int_{B_{R}}\frac{1}{t}\int_{0}^{t}|\psi_{exp}(r,\xi)|^{t\Re[\beta(r)]}\mathrm{d}r\mathcal{F}[u_{0}](\xi)\mathrm{d}\xi$
$\quad+\exp\left(2\pi\int_{0}^{t}|\Im[\beta(r)]|\mathrm{d}r\right)\int_{0}^{t}\int_{B_{R}}\frac{1}{t-s}\int_{s}^{t}\left(|\psi_{exp}(r,\xi)|^{(t-s)\Re[\beta(r)]}\right)\mathrm{d}r\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\xi.$
Thus the above inequalities with (2.16) and (2.17) show that (2.9) holds. Next, observe that
$|\psi(\rho,\xi)|\leq|\beta(\rho)|\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|+2\pi|\beta(\rho)|$
if one considers the standard Branch in the complex logarithm. Therefore for all $t\in(0,T)$ and $R\in(0,\infty)$,
$\int_{B_{R}}\int_{0}^{t}|\psi(\rho,\xi)|\exp\left(\int_{0}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\mathcal{F}[u_{0}](\xi)\mathrm{d}\rho\mathrm{d}\xi$
$+\int_{B_{R}}\int_{0}^{t}|\psi(\rho,\xi)|\int_{0}^{\rho}\exp\left(\int_{s}^{\rho}\Re[\psi(r,\xi)]\mathrm{d}r\right)\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\mathrm{d}\rho\mathrm{d}\xi$
$\lesssim\int_{B_{R}}\left[\int_{0}^{t}|\beta(\rho)|\left(1+\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|\right)\left(\frac{1}{\rho}\int_{0}^{\rho}|\psi_{exp}(r,\xi)|^{\rho\Re[\beta(r)]}\mathrm{d}r\right)\mathrm{d}\rho\mathcal{F}[u_{0}](\xi)\right]\mathrm{d}\xi$
$\quad+\int_{B_{R}}\Bigg{[}\int_{0}^{t}\left(\int_{s}^{t}|\beta(\rho)|\left(1+\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|\right)\left(\frac{1}{\rho-s}\int_{s}^{\rho}\left(|\psi_{exp}(r,\xi)|^{(\rho-s)\Re[\beta(r)]}\right)\mathrm{d}r\right)\mathrm{d}\rho\right)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left|\mathcal{F}[f(s,\cdot)](\xi)\right|\mathrm{d}s\Bigg{]}\mathrm{d}\xi,$ (6.1)
where
$\exp\left(2\pi\int_{0}^{t}|\Im[\beta(r)]|\mathrm{d}r\right)\leq\exp\left(2\pi\int_{0}^{t}|\beta(r)|\mathrm{d}r\right)<\infty$
is used in the inequality above. Finally, applying (2.18) to (6.1), we have (2.10). The theorem is proved. ∎
Next, we consider the special case of $\psi_{\exp}(t,\xi)$. More precisely, we consider the second-order case, i.e.
$\psi_{\exp}(t,\xi)=\alpha^{ij}(t)\xi^{i}\xi^{j}.$
Recalling
$\log\left(\psi_{exp}(t,-i\nabla)\right)u(t,x):=\mathcal{F}^{-1}\left[\log\left(\psi_{exp}(t,\xi)\right)\mathcal{F}[u(t,\cdot)](\xi)\right](x),$
we use the special notation
$\beta(t)\log\left(-\alpha^{ij}(t)D_{x^{i}}D_{x^{j}}\right)u(t,x)$
to denote
$\mathcal{F}^{-1}\left[\log\left(\alpha^{ij}(t)\xi^{i}\xi^{j}\right)\mathcal{F}[u(t,\cdot)](\xi)\right](x).$
Combining all the information above, we finally state an evolution equation with the operator $\beta(t)\log\left(-\alpha^{ij}(t)D_{x^{i}}D_{x^{j}}\right)$ as follows:
$\begin{cases}\partial_{t}u(t,x)=\beta(t)\log\left(-\alpha^{ij}(t)D_{x^{i}}D_{x^{j}}\right)u(t,x)+f(t,x),\quad&(t,x)\in(0,T)\times\mathbb{R}^{d},\\ u(0,x)=u_{0},\quad&x\in\mathbb{R}^{d}.\end{cases}$ (6.2)
We only consider real-valued symmetric $\left(\alpha^{ij}(t)\right)$. Additionally, assume that the coefficients $\alpha^{ij}(t)$ satisfy a uniform ellipticity, i.e. for each $t\in(0,T)$ there exists a positive constant $\nu$ such that
$\nu|\xi|^{2}\leq\alpha^{ij}(s)\xi^{i}\xi^{j}\leq\frac{1}{\nu}|\xi|^{2}\quad\forall(s,\xi)\in(0,t)\times\mathbb{R}^{d},$ (6.3)
[…]
Corollary 6.1.
Let $u_{0}\in\mathcal{F}^{-1}L_{1,\ell oc}(\mathbb{R}^{d})$ and $f\in\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)$. Assume that (6.3) holds. Additionally, suppose that
$0<\operatorname*{ess\,inf}_{s\in(0,t)}|\beta(s)|\leq\operatorname*{ess\,sup}_{s\in(0,t)}|\beta(s)|<\infty\quad\forall t\in(0,T).$ (6.4)
Then there exists a unique Fourier-space weak solution $u$ to (6.2) in
$\mathcal{F}^{-1}L_{\infty,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d}\right)\cap\mathcal{F}^{-1}L_{1,1,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},|\beta(t)|\mathrm{d}t\left|\log\left(\alpha^{ij}(t)\xi^{i}\xi^{j}\right)\right|\mathrm{d}\xi\right).$
Proof.
We use Theorem 2.27 to obtain this corollary. Let $t\in(0,T)$. Put
$\psi_{exp}(r,\xi)=\alpha^{ij}(r)\xi^{i}\xi^{j}\quad\forall r\in(0,t),$
$m=\operatorname*{ess\,inf}_{s\in(0,t)}|\beta(s)|,$
$M=\operatorname*{ess\,sup}_{s\in(0,t)}\left(|\beta(s)|+\sum_{i,j}|\alpha^{ij}(s)|\right).$
Due to (6.3) and (6.4), it is obvious that $m,M\in(0,\infty)$. It is sufficient to show that (2.17) and (2.18) hold. Applying (6.3) and (6.4), for all $t\in(0,T)$ and $R\in(0,\infty)$, we have
$\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\frac{1}{t-s}\int_{s}^{t}|\psi_{exp}(r,\xi)|^{(t-s)\Re[\beta(r)]}\mathrm{d}r\right)\lesssim\left(1+R^{2tM}\right)$
and
$\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\int_{s}^{t}|\beta(\rho)|\left(1+\left|\log\left(|\psi_{\exp}(\rho,\xi)|\right)\right|\right)\left(\frac{1}{\rho-s}\int_{s}^{\rho}\left(|\psi_{exp}(r,\xi)|^{(\rho-s)\Re[\beta(r)]}\right)\mathrm{d}r\right)\mathrm{d}\rho\right)$
$\lesssim\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(\int_{s}^{t}\left(1+\left|\log\left(|\xi|\right)\right|\right)\left(|\xi|^{(\rho-s)m}+|\xi|^{(\rho-s)M}\right)\mathrm{d}\rho\right)$
$\lesssim\sup_{(s,\xi)\in[0,t)\times B_{R}}\left(|\xi|^{(t-s)m}+|\xi|^{(t-s)M}\right)\lesssim\left(1+R^{tM}\right)<\infty.$
The corollary is proved. ∎
Remark 6.2.
The local boundedness of coefficients in (6.3) and (6.4) could be weakened by considering substitutes with complicated local integrabilities. It could be obtained by another application of Theorem 2.27. However, we do not give the details since conditions become extremely technical.
7. Second-order partial differential equations and proof of Theorem 2.34
We study (2.19) as a particular case of (1.5) in this section. We present a proof of Theorem 2.34 and give some applications of this theorem when data in weighted Bessel potential spaces. First, we prove Theorem 2.34.
Proof of Theorem 2.34.
Based on definitions, it is easy to check that (2.19) becomes a special case of (1.5) with the symbol
$\psi(t,\xi)=-a^{ij}(t)\xi^{i}\xi^{j}+ib^{j}(t)\xi^{j}+c(t).$
as mentioned in Remark 2.33. We use Corollary 5.3 to prove this theorem. Thus it is sufficient to show that
$\left\|\frac{1+|a^{ij}(\rho)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)|}{W_{k}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[a^{ij}(r)]\xi^{i}\xi^{j}+\Im[b^{j}(r)]\xi^{j}-\Re[c(r)]\right|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
is finite for all $t\in(0,T)$, $R\in(0,\infty)$, and $k=0,1$. It is an easy application of Minkowski’s and Hölder’s inequalities with (2.22), (2.23) and (2.24). Indeed, for each $k=0,1$, $t\in(0,T)$ and $R\in(0,\infty)$, applying (2.22), (2.23), and (2.24), we have
$\left\|\frac{1+|a^{ij}(\rho)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)|}{W_{k}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[a^{ij}(r)]\xi^{i}\xi^{j}+\Im[b^{j}(r)]\xi^{j}-\Re[c(r)]\right|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
$\lesssim\frac{1+\sum_{i,j}\|a^{ij}\|_{L_{p}\left((0,t)\right)}R^{2}+\sum_{j}\|b^{j}\|_{L_{p}\left((0,t)\right)}R+\|c\|_{L_{p}\left((0,t)\right)}}{\kappa_{k}(R)}$
×$\qquad\times\exp\left(\int_{0}^{t}\left[R^{2}\sum_{i,j}|a^{ij}(r)|+R\sum_{j}|b^{j}(r)|+|c(r)|\right]\mathrm{d}r\right)$
$\lesssim\frac{1+\sum_{i,j}\|a^{ij}\|_{L_{p}\left((0,t)\right)}R^{2}+\sum_{j}\|b^{j}\|_{L_{p}\left((0,t)\right)}R+\|c\|_{L_{p}\left((0,t)\right)}}{\kappa_{k}(R)}$
×$\qquad\times\exp\left(C\left(\sum_{i,j}\|a^{ij}\|_{L_{p}\left((0,t)\right)}R^{2}+\sum_{j}\|b^{j}\|_{L_{p}\left((0,t)\right)}R+\|c\|_{L_{p}\left((0,t)\right)}\right)\right)<\infty,$
where $C$ is a positive constant depending only on $t$ and $p$. The theorem is proved. ∎
We examine three distinct scenarios derived from Theorem 2.34, focusing on cases where the data are presented within the (weighted) Bessel potential spaces. We consider complex-valued coefficients so that $a^{ij},b^{j},c\in L_{p,loc}\left((0,T)\right)$ for all $i,j\in\{1,\ldots,d\}$ in the following corollaries.
Corollary 7.1.
Let $p\in[1,\infty]$, $q\in[1,2]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, $u_{0}\in H_{q}^{\gamma_{1}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\gamma_{2}}_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
$\mathcal{F}^{-1}L_{p,q^{\prime},t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t\left(1+\left|a^{ij}(t)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)\right|\right)\mathrm{d}\xi\right).$
Corollary 7.2.
Let $p\in[1,\infty]$, $q\in(1,2]$, $\gamma_{1},\tilde{\gamma}_{1}\in\left(-\frac{d(p-1)}{p},\infty\right)$, $\gamma_{2},\tilde{\gamma}_{2}\in[0,\infty)$, $u_{0}\in H_{q,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
$\mathcal{F}^{-1}L_{p,q^{\prime},t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t\left(1+\left|a^{ij}(t)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)\right|\right)\mathrm{d}\xi\right).$
Corollary 7.3.
Let $p\in[1,\infty]$, $q\in(2,\infty]$, $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $\gamma_{2},\tilde{\gamma}_{2}\in\left(\frac{d(q-2)}{2(q-1)},\infty\right)$, $u_{0}\in H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
$\mathcal{F}^{-1}L_{p,2,t\text{-}loc,x\text{-}\ell oc}\left((0,T)\times\mathbb{R}^{d},\mathrm{d}t\left(1+\left|a^{ij}(t)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)\right|\right)\mathrm{d}\xi\right).$
The proofs of these corollaries could be easily obtained from Theorem 2.34 with some properties of the weighted Bessel potential spaces developed in Section 3. For the details, follow the proofs of Corollaries 5.4, 5.6, and 5.7.